### v2

# On moduli spaces of canonical threefolds with small genera and minimal volumes

Abstract.
We prove that the canonical model of a $3$-fold of general type with geometric genus $2$ and with minimal canonical volume $\frac{1}{3}$ must be a hypersurface of degree $16$ in $\mathbb{P}(1,1,2,3,8)$, which gives an explicit description of its canonical ring. This implies that the coarse moduli space $\mathcal{M}_{\frac{1}{3},2}$, parametrizing all canonical $3$-folds with canonical volume $\frac{1}{3}$ and geometric genus $2$, is an irreducible unirational variety of dimension $189$. Parallel studies show that $\mathcal{M}_{1,3}$ is irreducible unirational as well and is of dimension $236$, and that $\mathcal{M}_{2,4}$ is irreducible unirational and is of dimension $270$. As being conceived, every member in these 3 families is simply connected. Additionally, our method yields $\textrm{Vol}\geq\frac{4}{3}p_{g}-\frac{10}{3}$ for $3$-folds of general type with $5\leq p_{g}\leq 10$, which completely solves all remaining cases of the Noether inequality of $3$-folds.
Key words and phrases:
canonical threefold, moduli space, canonical volume, weighted hypersurface
2020 Mathematics Subject Classification:
14J30, 14J10, 14D22
1. Introduction
Within birational geometry, perhaps the last phase of classifying projective varieties is to construct and to study the geometry of the concrete moduli space for each designated class. This is corroborated by the theory for curves and surfaces.
In higher dimensions, the birational classification of smooth projective varieties of general type splits into two steps: the first step is to find a good representative in each birational equivalence class, namely, the canonical model, and this step is completed by the minimal model program (see [21, 25, 1]); the second step is to study the geometry of canonical models and the moduli space of canonical models with given discrete numerical invariants (see [26, 33, 24]). While the general moduli theory is well-established, the explicit geometry of certain concrete moduli spaces is not yet well-understood.
This paper aims to investigate the birational geometry and moduli spaces of several special families, of $3$-folds of general type, with small genera and attaining minimal canonical volumes.
We start by recalling the following theorem:
Theorem 1 ([7, Theorem 1.4, Theorem 1.5]).
Let $W$ be a smooth projective $3$-fold of general type. Denote by $\text{\rm Vol}(W)$ the canonical volume of $W$. Then
  1. (1)
    when $p_{g}(W)\geq 2$, $\text{\rm Vol}(W)\geq\frac{1}{3}$;
  2. (2)
    when $p_{g}(W)\geq 3$, $\text{\rm Vol}(W)\geq 1$;
  3. (3)
    when $p_{g}(W)\geq 4$, $\text{\rm Vol}(W)\geq 2$.
Furthermore, the above three lower bounds $\frac{1}{3}$, $1$ and $2$ are all optimal and the minimal volume is attained only if $p_{g}=2,3,4$, respectively.
It is a natural question to study the classification problem for those $3$-folds attaining the above minimal volumes, as analogue to curves with small genus or surfaces with small $c_{1}^{2}$. Also, it is expected that the classification of such $3$-folds with extreme numerical values will benefit the development of explicit classification theory of higher dimensional algebraic varieties.
To study the birational geometry of a variety $W$ of general type, one classical method is to study its canonical ring $R(W):=\bigoplus_{k\geq 0}H^{0}(W,kK_{W})$ which originates from the work of Zariski and Mumford ([34]). Reid proposed a program to study the canonical ring from the point of view of commutative algebra which predicts the construction of $W$ (as a subvariety in a weighted projective space) from the algebraic structure of $R(W)$, see [29, 31]. However, to the best of our knowledge, the explicit algebraic structure of the canonical ring of a variety is only known in the case that the given variety has a very clear geometric structure (e.g. a complete intersection in a weighted projective space). In this paper, we will give explicit descriptions for the canonical rings of smooth projective $3$-folds of general type with $2\leq p_{g}(X)\leq 4$ attaining minimal volumes.
For a positive rational number $a$ and a nonnegative integer $b$, denote by $\mathcal{M}_{a,b}$ the coarse moduli space parametrizing all projective $3$-folds $X$ with canonical singularities such that $K_{X}$ is ample, $K_{X}^{3}=a$, and $p_{g}(X)=b$ in the sense of Viehweg [33, §8.5] and Kollár–Shepherd-Barron [26, §5.4] (see also [24]). When $\mathcal{M}_{a,b}$ is nonempty, one would like to know its geometric properties as detailed as possible, for example, the number of its irreducible or connected components and the dimension of each component. A further task is to construct an ideal compactification for ${\mathcal{M}_{a,b}}$, and one successful way is the KSB moduli theory via the minimal model program (see [26, 24]). To the best of our knowledge, for the explicit geometry of moduli spaces of $3$-folds of general type, very few is known in the literature. In this paper, we mainly study the irreducibility and dimensions of $\mathcal{M}_{\frac{1}{3},2}$, $\mathcal{M}_{1,3}$ and $\mathcal{M}_{2,4}$. It is not out of expectation that every member in these families will be proven to be simply connected.
Our main results are as follows.
Theorem 1.1.
Let $W$ be a smooth projective $3$-fold of general type with $p_{g}(W)\geq 2$ and $\text{\rm Vol}(W)=\frac{1}{3}$. Then
  1. (1)
    $W$ has the same plurigenera as those of a general hypersurface of degree $16$ in $\mathbb{P}(1,1,2,3,8)$;
  2. (2)
    the canonical model of $W$ is a hypersurface in $\mathbb{P}(1,1,2,3,8)$ defined by a weighted homogeneous polynomial $f$ of degree $16$, where
    $f(x_{0},\dots,x_{4})=x_{4}^{2}+f_{0}(x_{0},x_{1},x_{2},x_{3})$
    in suitable homogeneous coordinates $[x_{0}:\dots:x_{4}]$ of $\mathbb{P}(1,1,2,3,8)$; in other words, the canonical ring of $W$ is
    $R(W)\simeq\mathbb{C}[x_{0},\dots,x_{4}]/(f),$
    where $\mathbb{C}[x_{0},\dots,x_{4}]$ is viewed as a weighted polynomial ring with $\operatorname{wt}(x_{0},\dots,x_{4})=(1,1,2,3,8)$;
  3. (3)
    $W$ is simply connected;
  4. (4)
    the moduli space $\mathcal{M}_{\frac{1}{3},2}$ is irreducible, unirational, and $\dim\mathcal{M}_{\frac{1}{3},2}=189$.
Theorem 1.2.
Let $W$ be a smooth projective $3$-fold of general type with $p_{g}(W)\geq 3$ and $\text{\rm Vol}(W)=1$. Then
  1. (1)
    $W$ has the same plurigenera as those of a general hypersurface of degree $12$ in $\mathbb{P}(1,1,1,2,6)$;
  2. (2)
    the canonical model of $W$ is a hypersurface in $\mathbb{P}(1,1,1,2,6)$ defined by a weighted homogeneous polynomial $f$ of degree $12$, where
    $f(x_{0},\dots,x_{4})=x_{4}^{2}+f_{0}(x_{0},x_{1},x_{2},x_{3})$
    in suitable homogeneous coordinates $[x_{0}:\dots:x_{4}]$ of $\mathbb{P}(1,1,1,2,6)$; in other words, the canonical ring of $W$ is
    $R(W)\simeq\mathbb{C}[x_{0},\dots,x_{4}]/(f),$
    where $\mathbb{C}[x_{0},\dots,x_{4}]$ is viewed as a weighted polynomial ring with $\operatorname{wt}(x_{0},\dots,x_{4})=(1,1,1,2,6)$;
  3. (3)
    $W$ is simply connected;
  4. (4)
    the moduli space $\mathcal{M}_{1,3}$ is irreducible, unirational, and $\dim\mathcal{M}_{1,3}=236$.
Theorem 1.3.
Let $W$ be a smooth projective $3$-fold of general type with $p_{g}(W)\geq 4$ and $\text{\rm Vol}(W)=2$. Then
  1. (1)
    $W$ has the same plurigenera as those of a general hypersurface of degree $10$ in $\mathbb{P}(1,1,1,1,5)$;
  2. (2)
    the canonical model of $W$ belongs to either of the following $2$ types:
    1. (a)
      Type I: a hypersurface in $\mathbb{P}(1,1,1,1,5)$ defined by a weighted homogeneous polynomial $f$ of degree $10$, where
      $f(x_{0},\dots,x_{4})=x_{4}^{2}+f_{0}(x_{0},x_{1},x_{2},x_{3})$
      in suitable homogeneous coordinates $[x_{0}:\dots:x_{4}]$ of $\mathbb{P}(1,1,1,1,5)$; in other words, the canonical ring of $W$ is
      $R(W)\simeq\mathbb{C}[x_{0},\dots,x_{4}]/(f),$
      where $\mathbb{C}[x_{0},\dots,x_{4}]$ is viewed as a weighted polynomial ring with $\operatorname{wt}(x_{0},\dots,x_{4})=(1,1,1,1,5)$;
    2. (b)
      Type II: a subvariety in $\mathbb{P}(1,1,1,1,2,5)$ defined by $2$ weighted homogeneous polynomials $q$ and $f$ of degrees $2$ and $10$ respectively, where
      $q(x_{0},x_{1},x_{2},x_{3}){}$ $=x_{3}^{2}+q_{0}(x_{0},x_{1},x_{2}),$
      $f(x_{0},\dots,x_{5}){}$ $=x_{5}^{2}+f_{0}(x_{0},\dots,x_{4})$
      in suitable homogeneous coordinates $[x_{0}:\dots:x_{5}]$ of $\mathbb{P}(1,1,1,1,2,5)$; in other words, the canonical ring of $W$ is
      $R(W)\simeq\mathbb{C}[x_{0},\dots,x_{5}]/(q,f),$
      where $\mathbb{C}[x_{0},\dots,x_{5}]$ is viewed as a weighted polynomial ring with $\operatorname{wt}(x_{0},\dots,x_{5})=(1,1,1,1,2,5)$;
    furthermore, a canonical $3$-fold of type II is a specialization of canonical $3$-folds of type I;
  3. (3)
    $W$ is simply connected;
  4. (4)
    the moduli space $\mathcal{M}_{2,4}$ is irreducible, unirational, and $\dim\mathcal{M}_{2,4}=270$.
As a by-product of this paper, we obtain the following result which completely solves all remaining cases ($5\leq p_{g}\leq 10$) of the Noether inequality in the previous work of Chen–Chen–Jiang [4, 5].
Theorem 1.4 (=Theorem 9.5).
Let $X$ be a minimal $3$-fold of general type with $5\leq p_{g}(X)\leq 10$. Then the Noether inequality $K_{X}^{3}\geq\frac{4}{3}p_{g}(X)-\frac{10}{3}$ holds.
We briefly explain the main methods and difficulties of this paper. Consider Theorem 1.1 for instance, which is the most difficult one. The key part is to show that the canonical model can be embedded into $\mathbb{P}(1,1,2,3,8)$. As in [19, 20], we have an effective criterion on how to embed a given polarized variety into a certain weighted projective space. In the context of Theorem 1.1, to apply this criterion, we need to know all the plurigenera of $W$ and the behavior of pluricanonical systems of $W$. Usually plurigenera can be computed by Reid’s Riemann–Roch formula on a minimal model $X$ of $W$, however, in order to apply this formula, only knowing the geometric genus and the canonical volume is not sufficient, one has to know the singularities of $X$, as well as the holomorphic Euler characteristic of $X$. We put our main effort to get such geometric information.
In the study of explicit geometry of $3$-folds of general type, one of the most challenging cases is when $X$ admits a pencil $|F_{X}|$ which is not free. One key observation (see §3) is, by applying MMP to a log resolution of the pair $(X,F_{X})$ in question, to obtain a relatively minimal model $(X^{\prime},F)$ over $X$ on which $F$ is free. Such a birational model can be viewed as a minimal partial resolution for the pencil $|F_{X}|$. With the help of this new birational model, we can efficiently analyze the geometry of $X$, and get the information on the singularities of $X$. Then we analyze the behavior of pluricanonical systems of $X$ which gives restrictions on the plurigenera of $X$. Together with Reid’s Riemann–Roch formula, such restrictions will eventually lead us to the precise value of the holomorphic Euler characteristic of $X$ and all plurigenera of $X$.
This paper is organized as the following. In §2, we recall some basic knowledge and background. In §3, we construct a relative minimal model resolving a pencil by the minimal model program. In §4–§6, we study the geometry, of minimal $3$-folds of general type with $2\leq p_{g}\leq 4$ and with minimal canonical volume, such as their singularities, pluricanonical systems and plurigenera. In §7, we recall the criterion in [20] in a slightly more general form. In §8, we prove our main results upon the preparation of previous sections. In §9, we prove the Noether inequality for $5\leq p_{g}\leq 10$ as an application of the birational model in §3.
2. Preliminaries
Throughout this paper, we work over the complex number field $\mathbb{C}$ meanwhile all results can be easily generalized to any algebraically closed field of characteristic 00.
We adopt the standard notation and definitions in [21, 25], and will freely use them.
A ${\mathbb{Q}}$-divisor is said to be ${\mathbb{Q}}$-effective if it is ${\mathbb{Q}}$-linearly equivalent to an effective ${\mathbb{Q}}$-divisor.
2.1. Singularities
A log pair $(X,B)$ consists of a normal projective variety $X$ and an effective $\mathbb{Q}$-divisor $B$ on $X$ such that $K_{X}+B$ is $\mathbb{Q}$-Cartier.
Let $f:Y\to X$ be a log resolution of the log pair $(X,B)$, write
$K_{Y}=f^{*}(K_{X}+B)+\sum a_{i}D_{i},$
where $\{D_{i}\}$ are distinct prime divisors and the sum runs through all irreducible components in the union of $f$-exceptional divisors and the strict transform of $B$. The log pair $(X,B)$ is called
  1. (a)
    kawamata log terminal (klt, for short) if $a_{i}>-1$ for all $i$;
  2. (b)
    purely log terminal (plt, for short) if $a_{i}>-1$ for all $f$-exceptional divisors $F_{i}$ and all $f$;
  3. (c)
    terminal if $a_{i}>0$ for all $f$-exceptional divisors $D_{i}$ and all $f$;
  4. (d)
    canonical if $a_{i}\geq 0$ for all $f$-exceptional divisors $D_{i}$ and all $f$.
Usually we write $X$ instead of $(X,0)$ in the case $B=0$.
2.2. Minimal model and canonical model
A normal projective variety $X$ is said to be minimal if $X$ has $\mathbb{Q}$-factorial terminal singularities and $K_{X}$ is nef.
A normal projective variety $X$ with canonical singularities is called of general type if $K_{X}$ is big.
According to the minimal model program (see [1]), a smooth projective variety $W$ of general type is always birational to a minimal model $W_{\min}$ which is a minimal variety of general type, and to a unique canonical model $W_{\textrm{can}}$ which is a normal projective variety with canonical singularities and $K_{W_{\textrm{can}}}$ is ample. In this case, the canonical volume of $W$ is defined to be the self-intersection number of the canonical divisor of a minimal model or the canonical model, i.e., $\text{\rm Vol}(W):=K_{W_{\min}}^{\dim W}=K_{W_{\textrm{can}}}^{\dim W}$. The geometric genus of $W$ is defined by $p_{g}(W):=h^{0}(W,K_{W})=h^{0}(W_{\min},K_{W_{\min}}).$
In the study of birational geometry of a smooth projective variety of general type, we often freely replace it by its minimal model or canonical model.
For two integers $m>0$ and $n\geq 0$, an $(m,n)$-surface is a smooth projective surface $S$ of general type with $\text{\rm Vol}(S)=m$ and $p_{g}(S)=n$.
2.3. Rational maps defined by linear systems
Let $X$ be a normal projective variety on which $D$ is a ${\mathbb{Q}}$-Cartier Weil divisor with $h^{0}(X,D)\geq 2$. Take a linear subspace $V\subseteq H^{0}(X,D)$ with $\dim V\geq 2$ and consider the corresponding sublinear system $\Lambda\subseteq|D|$. We can consider the rational map defined by $\Lambda$, say $\Phi_{\Lambda}:X{\dashrightarrow}{\mathbb{P}}^{\dim V-1}$, which is not necessarily well-defined everywhere. By Hironaka’s big theorem, we can take successive blow-ups $\pi:W\to X$ such that:
  1. (i)
    $W$ is smooth and projective;
  2. (ii)
    corresponding to $\Lambda$, the sublinear system $\Lambda_{W}\subseteq|\lfloor{\pi^{*}(D)}\rfloor|$ has free movable part $\operatorname{Mov}(\Lambda_{W})$ and, consequently, the rational map $\gamma=\Phi_{\Lambda}\circ\pi$ is a morphism.
[…]
${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{\gamma}$$\scriptstyle{\psi}$${\Sigma^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\Lambda}}$${\Sigma}$
  1. (1)
    If $\dim(\Sigma^{\prime})\geq 2$, then a general member of $\operatorname{Mov}(\Lambda_{W})$ is a smooth projective variety by Bertini’s theorem, and is called a general irreducible element of $\operatorname{Mov}(\Lambda_{W})$ (or $\Lambda_{W}$).
  2. (2)
    If $\dim(\Sigma^{\prime})=1$, then a general fiber $S$ of $\psi$ is a smooth projective variety by Bertini’s theorem. In this case, we say that $\Lambda$ is composed with a pencil and $S$ is called a general irreducible element of $\operatorname{Mov}(\Lambda_{W})$ (or $\Lambda_{W}$). $\Lambda$ is said to be rational if $\Sigma^{\prime}\simeq{\mathbb{P}}^{1}$. A general member of $\operatorname{Mov}(\Lambda_{W})$ is of the form $\sum_{i=1}^{a}S_{i}$, where $S_{i}$ is a smooth fiber of $\psi$ for each $i$. It is clear that $a=\dim V-1$ if $\Sigma^{\prime}\simeq{\mathbb{P}}^{1}$, and $a\geq\dim V$ if $\Sigma^{\prime}\not\simeq{\mathbb{P}}^{1}$.
We will frequently use the following setting.
Setting 2.1.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)\geq 2$. Take a birational modification $\pi:W\to X$ as in §2.3 with respect to the canonical system $|K_{X}|$ and keep the same notation. Write
$K_{W}{}$ $=\pi^{*}(K_{X})+E_{\pi},$
$\pi^{*}(K_{X}){}$ $=M+Z_{\pi},$
where $|M|=\operatorname{Mov}|\lfloor{\pi^{*}(K_{X})}\rfloor|$, $E_{\pi}$ and $Z_{\pi}$ are effective $\mathbb{Q}$-divisors. Let $S$ be a general irreducible element of $|M|$. After taking some further blow-up, we may further assume that the following hold:
  1. (1)
    $\operatorname{Mov}|K_{S}|$ is base point free;
  2. (2)
    for any integers $0<m\leq 4$ and $0\leq j\leq m$, the linear system
    $|M_{m,-j}|:=\operatorname{Mov}|\lfloor{\pi^{*}(mK_{X})-jS}\rfloor|$
    is base point free or empty;
  3. (3)
    $\operatorname{Supp}(Z_{\pi}+E_{\pi})$ is of simple normal crossing.
Consider the natural restriction map
$\theta_{m,-j}:H^{0}(W,M_{m,-j})\to H^{0}(S,M_{m,-j}|_{S}),$
and set $u_{m,-j}=\dim\operatorname{Im}\theta_{m,-j}$.
For two linear systems $\Lambda_{1}$ and $\Lambda_{2}$, we write $\Lambda_{1}\succcurlyeq\Lambda_{2}$ if there exists an effective divisor $N$ such that $\Lambda_{1}\supseteq\Lambda_{2}+N$.
2.4. Reid’s Riemann–Roch formula
A basket is a collection of pairs of integers (permitting weights).
Let $X$ be a projective terminal $3$-fold. According to Reid [30], there is a basket of orbifold points (called the Reid basket)
$B_{X}=\bigg{\{}(b_{i},r_{i})\mid i=1,\cdots,s;0<b_{i}\leq\frac{r_{i}}{2};b_{i}\text{ is coprime to }r_{i}\bigg{\}}$
associated to $X$, which comes from locally deforming the singularities of $X$ into cyclic quotient singularities, where a pair $(b_{i},r_{i})$ corresponds to a (virtual) orbifold point $Q_{i}$ of type $\frac{1}{r_{i}}(1,-1,b_{i})$.
By Reid’s Riemann–Roch formula [30, Corollary 10.3], for any positive integer $m$,
$\chi(X,{\mathcal{O}}_{X}(mK_{X}))$
[…]
where $l(m)=\sum_{i}\sum_{j=1}^{m-1}\frac{\overline{jb_{i}}(r_{i}-\overline{jb_{i}})}{2r_{i}}$ and the first sum runs over all orbifold points in the Reid basket. Here $\overline{jb_{i}}$ means the smallest nonnegative residue of $jb_{i}\bmod r_{i}$.
For a positive integer $m$, the $m$-th plurigenus of $X$ is denoted by $P_{m}(X):=h^{0}(X,mK_{X})$, and it can be computed by Reid’s Riemann–Roch formula and the Kawamata–Viehweg vanishing theorem when $X$ is minimal of general type and $m\geq 2$.
2.5. Hodge index theorem
We recall the following well-known result which is a consequence of the Hodge index theorem.
Lemma 2.2.
Let $D_{1},D_{2}$ be $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisors on a normal projective surface $S$.
  1. (1)
    If $D_{1},D_{2}$ are nef, $D_{1}^{2}=D_{2}^{2}>0$, and $D_{1}-D_{2}$ is effective, then $D_{1}=D_{2}$.
  2. (2)
    If $D_{1},D_{2}$ are effective, $(D_{1}\cdot D_{2})=0$, and $D_{1}+D_{2}$ is nef and big, then either $D_{1}=0$ or $D_{2}=0$.
Proof.
Recall that the Hodge index theorem states that if $H_{1}$ and $H_{2}$ are $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisors on $S$ such that $H_{1}^{2}>0$ and $(H_{1}\cdot H_{2})=0$, then $H_{2}^{2}\leq 0$; moreover, the equality holds if and only if $H_{2}\equiv 0$. Here recall that if $H_{2}$ is effective, then $H_{2}\equiv 0$ if and only if $H_{2}=0$.
(1) Denote $D=D_{1}-D_{2}\geq 0$. Since $0=D_{1}^{2}-D_{2}^{2}=((D_{1}+D_{2})\cdot D)$ and $D_{1},D_{2}$ are nef, we have $(D_{1}\cdot D)=(D_{2}\cdot D)=0$. In particular, $D^{2}=0$. On the other hand, $(D_{1}+D_{2})^{2}>0$. So by the Hodge index theorem, $D=0$.
(2) Note that $D_{1}^{2}+D_{2}^{2}=(D_{1}+D_{2})^{2}>0$. So either $D_{1}^{2}>0$ or $D_{2}^{2}>0$. Without loss of generality, we may assume that $D_{1}^{2}>0$. Then $(D_{1}\cdot D_{2})=0$ implies that $D_{2}^{2}\leq 0$ by the Hodge index theorem. On the other hand, $D_{2}^{2}=(D_{2}\cdot(D_{1}+D_{2}))\geq 0$ as $D_{1}+D_{2}$ is nef. Hence $D_{2}^{2}=0$ and $D_{2}=0$ by the Hodge index theorem. ∎
3. A relatively minimal model resolving a pencil
In this section, we construct a birational model which resolves a given pencil by using the minimal model program.
Proposition 3.1.
Let $X$ be a normal $\mathbb{Q}$-factorial projective variety on which a linear system $\Lambda$ is composed with a pencil. Then there exists a projective variety $X^{\prime}$ with a birational morphism $g:X^{\prime}\to X$ having the following properties:
  1. (1)
    $X^{\prime}$ is $\mathbb{Q}$-factorial terminal;
  2. (2)
    the movable part of the linear system $\Lambda^{\prime}$, on $X^{\prime}$, corresponding to $\Lambda$ is base point free and, after taking the Stein factorization, ${\Lambda^{\prime}}$ induces a fibration $f:X^{\prime}\to\Gamma$ onto a curve $\Gamma$;
  3. (3)
    for a general fiber $F$ of $f$, $K_{X^{\prime}}+F$ is $g$-nef;
  4. (4)
    $g^{*}(K_{X}+g_{*}F)-K_{X^{\prime}}-F$ is an effective $g$-exceptional $\mathbb{Q}$-divisor.
Proof.
Take a birational modification $\pi:W\to X$ and a fibration $\psi:W\to\Gamma$ as in §2.3 induced by $\Lambda$. Take $S_{1},S_{2}$ to be two general fibers of $\psi$, then $S_{1}$ and $S_{2}$ are disjoint and smooth. The pair $(W,\frac{1}{2}S_{1}+\frac{1}{2}S_{2})$ is klt and terminal. By [1, Corollary 1.4.2], we may run a $(K_{W}+\frac{1}{2}S_{1}+\frac{1}{2}S_{2})$-MMP over $X$ which terminates at a model $g:X^{\prime}\to X$ such that $K_{X^{\prime}}+\frac{1}{2}F_{1}+\frac{1}{2}F_{2}$ is nef over $X$, where $F_{1}$ and $F_{2}$ are strict transforms of $S_{1}$ and $S_{2}$ on $X^{\prime}$. We claim that $X^{\prime}$ has all required properties.
As $S_{1}$ and $S_{2}$ are movable, they are not contracted by the MMP, so $(X^{\prime},\frac{1}{2}F_{1}+\frac{1}{2}F_{2})$ is terminal by [25, Corollary 3.42, Corollary 3.43]. In particular, $X^{\prime}$ is $\mathbb{Q}$-factorial terminal.
We claim that $F_{1}\cap F_{2}=\emptyset$. Suppose that $F_{1}\cap F_{2}\neq\emptyset$, then there exists a subvariety $Q\subseteq F_{1}\cap F_{2}$ of codimension $2$ in $X^{\prime}$ as $X^{\prime}$ is $\mathbb{Q}$-factorial. On the other hand, $X^{\prime}$ is smooth at the generic point of $Q$ by [25, Corollary 5.18]. So by [25, Lemma 2.29], the blowing up along $Q$ induces an exceptional prime divisor $E$ with discrepancy $a(E,X^{\prime},\frac{1}{2}F_{1}+\frac{1}{2}F_{2})\leq 1-\frac{1}{2}-\frac{1}{2}=0,$ but this contradicts the fact that $(X^{\prime},\frac{1}{2}F_{1}+\frac{1}{2}F_{2})$ is terminal. So $F_{1}\cap F_{2}=\emptyset$.
If $\Gamma\simeq\mathbb{P}^{1}$, then the free linear system $|F_{1}|$ induces a natural morphism $f:X^{\prime}\to\Gamma$. On the other hand, if $g(\Gamma)>0$, then all steps of the MMP are relative over $\Gamma$ since negative extremal rays are generated by rational curves, and hence $W\to\Gamma$ descends to a natural morphism $f:X^{\prime}\to\Gamma$. Clearly in either case, this morphism is defined by the linear system $\Lambda^{\prime}$, on $X^{\prime}$, corresponding to $\Lambda$.
Finally consider
$G=g^{*}(K_{X}+g_{*}F)-K_{X^{\prime}}-F.$
Then $G$ is exceptional over $X$ and
$-G\equiv_{X}K_{X^{\prime}}+F\equiv K_{X^{\prime}}+\frac{1}{2}F_{1}+\frac{1}{2}F_{2}$
is nef over $X$ by our construction. Hence $G\geq 0$ by the negativity lemma [25, Lemma 3.39]. ∎
As a consequence of Proposition 3.1, we will frequently use the following modification for a minimal $3$-fold of general type with nontrivial canonical map.
Setting 3.2.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)\geq 2$. Take a general sublinear system $\Lambda\subseteq|K_{X}|$ which is composed with a pencil. Applying Proposition 3.1 to $\Lambda$, we get a projective birational morphism $\mu:X^{\prime}\to X$ with a surjective fibration $f:X^{\prime}\to\Gamma$ such that $X^{\prime}$ is $\mathbb{Q}$-factorial terminal and
$\mu^{*}(K_{X}+F_{X})-K_{X^{\prime}}-F=G$ (3.1)
is an effective $\mu$-exceptional $\mathbb{Q}$-divisor, where $F$ is a general fiber of $f$ and $F_{X}=\mu_{*}F$. Here $G$ is independent of $F$ by the negativity lemma [25, Lemma 3.39]. We may write
$K_{X}{}$ $=F_{X}+Z_{X},$
$\mu^{*}(K_{X}{})$ $=F+Z,$
$K_{X^{\prime}}{}$ $=\mu^{*}(K_{X})+E_{\mu},$
where $Z_{X}$, $E_{\mu}$ and $Z$ are effective $\mathbb{Q}$-divisors and $E_{\mu}$ is $\mu$-exceptional. Note that $F$ is a smooth projective surface of general type. Denote by $\sigma:F\to F_{0}$ the contraction to its minimal model.
Lemma 3.3.
Keep the notation in Setting 3.2. Suppose that $|K_{X}|$ is composed with a rational pencil, then
$p_{g}(X)\mu^{*}(K_{X})|_{F}-(p_{g}(X)-1)\sigma^{*}(K_{F_{0}})$
is effective. More precisely, there exists an effective divisor $D\in|K_{X}|$ and an effective divisor $D_{0}\in|K_{F_{0}}|$ such that
$p_{g}(X)\mu^{*}(D)|_{F}-(p_{g}(X)-1)\sigma^{*}(D_{0})$
is an effective $\mathbb{Q}$-divisor.
Proof.
To simplify the notation we denote $d=p_{g}(X)-1$. Clearly $\Lambda$ and $|K_{X}|$ are composed with the same pencil. So we may write $K_{X}=dF_{X}+\Delta$, where $\Delta$ is the fixed part of $|K_{X}|$. Choose a general fiber $F^{\prime}$ of $f$ different from $F$, denote $F^{\prime}_{X}=\mu_{*}F^{\prime}$. Then we can take $D=dF^{\prime}_{X}+\Delta$.
We first explain the construction of $D_{0}$. As $D\in|K_{X}|$, we have $\mu^{*}(D)+E_{\mu}\in|K_{X^{\prime}}|$. Then $\mu^{*}(D)|_{F}+E_{\mu}|_{F}\in|K_{F}|$, where the restriction is well-defined as $F\not\subset\operatorname{Supp}(\mu^{*}(D)+E_{\mu})$. Then we just take $D_{0}:=\sigma_{*}(\mu^{*}(D)|_{F}+E_{\mu}|_{F})\in|K_{F_{0}}|$. Clearly $\mu^{*}(D)|_{F}+E_{\mu}|_{F}-\sigma^{*}(D_{0})=K_{F}-\sigma^{*}(K_{F_{0}})$ is effective.
Applying (3.1) to $D\in|K_{X}|$ and $F^{\prime}$, one has
$(d+1)\mu^{*}(D)\geq d\mu^{*}(D+F^{\prime}_{X})=d(\mu^{*}(D)+E_{\mu}+F^{\prime}+G)\geq d(\mu^{*}(D)+E_{\mu}).$
Then we get the conclusion by restricting on $F$. ∎
Lemma 3.4.
Keep the notation in Setting 3.2. Suppose that $(F,G|_{F})$ is klt.
  1. (1)
    Then $(X,F_{X})$ is plt, $F_{X}$ is normal and klt, and
    $K_{F}+G|_{F}=(\mu|_{F})^{*}(K_{F_{X}}).$
  2. (2)
    Suppose that $Z_{X}=0$. For a non-Gorenstein singularity $P$ of $X$, denote by $r_{P}$ the Cartier index of $K_{X}$ at $P$. Then $P\in F_{X}$ and $r_{P}$ equals to the Cartier index of $K_{X}|_{F_{X}}$ (as a well-defined Weil divisor on $F_{X}$) at $P$. Furthermore, if $P$ is a Gorenstein point on $F_{X}$, then $r_{P}=2$.
Proof.
(1) Recall that by (3.1),
$K_{X^{\prime}}+F+G=\mu^{*}(K_{X}+F_{X}).$
Since $({F},G|_{F})$ is klt, by [25, Theorem 5.50], $(X^{\prime},F+G)$ is plt in a neighborhood of $F$. Then by [25, Theorem 5.48], $(X,F_{X})$ is plt in a neighborhood of $F_{X}$. This implies that $(X,F_{X})$ is plt as $X$ is terminal. In particular, $F_{X}$ is normal by [25, Proposition 5.51]. The remaining assertions follow from the adjunction formula $(K_{X}+F_{X})|_{F_{X}}=K_{F_{X}}$ and [25, Theorem 5.50] as $X$ has only isolated singularities.
(2) Since $K_{X}\sim F_{X}$, all non-Gorenstein singularities of $X$ are contained in $F_{X}$. Recall that $K_{X}|_{F_{X}}$ is a well-defined Weil divisor on $F_{X}$ since $X$ and $F_{X}$ both have only isolated singularities and $F_{X}$ is general. Then $r_{P}$ equals to the Cartier index of $K_{X}|_{F_{X}}$ at $P$ by [14, Theorem A.1]. For the last assertion, note that $2K_{X}|_{F_{X}}\sim(K_{X}+F_{X})|_{F_{X}}=K_{F_{X}}$ by the adjunction formula. So if $K_{F_{X}}$ is Cartier at $P$, then $r_{P}=2$. ∎
Lemma 3.5.
Keep the notation in Setting 3.2. Then $Z_{X}=0$ if and only if $\mu^{*}(Z_{X})|_{F}\equiv 0$.
Proof.
We only need to show the “if” part. Take a general very ample divisor $H$ on $X$, by the projection formula,
$(F_{X}|_{H}\cdot Z_{X}|_{H})=(H\cdot Z_{X}\cdot F_{X})=(\mu^{*}(H)\cdot\mu^{*}(Z_{X})\cdot F)=0.$
On the other hand, $K_{X}|_{H}=F_{X}|_{H}+Z_{X}|_{H}$ is a nef and big divisor on $H$. Clearly $F_{X}|_{H}\neq 0$, so $Z_{X}|_{H}=0$ by Lemma 2.2(2), which implies that $Z_{X}=0$. ∎
4. Birational geometry of $3$-folds with $p_{g}=2$ and $\text{\rm Vol}=1/3$
In this section, we study the birational geometry of a minimal $3$-fold $X$ of general type with $p_{g}(X)=2$ and $K_{X}^{3}=\frac{1}{3}.$
4.1. Geometric structure and non-Gorenstein singularities
Since $p_{g}(X)=2$, $|K_{X}|$ is composed with a pencil. Take a birational modification $\mu:X^{\prime}\to X$ as in Setting 3.2 and keep the same notation. Since $h^{1}(\mathcal{O}_{X})=0$ by [7, Theorem 1.5(3)], we have $\Gamma\simeq\mathbb{P}^{1}$, in other words, $|K_{X}|$ is composed with a rational pencil. By [7, Theorem 1.4(2)], $F$ is a $(1,2)$-surface.
We recall the following lemma from [7, §2.15].
Lemma 4.1.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=2$ and $K_{X}^{3}=\frac{1}{3}$. Keep the notation in Setting 3.2. Then
$(\mu^{*}(K_{X})|_{F})^{2}=\frac{1}{3},\quad(\mu^{*}(K_{X})|_{F}\cdot\sigma^{*}(K_{F_{0}}))=(\mu^{*}(K_{X})|_{F}\cdot\operatorname{Mov}|\sigma^{*}(K_{F_{0}})|)=\frac{2}{3}.$
Proof.
For the convenience of readers who are not familiar with [7], we briefly recall the proof. Denote by $\xi=(\mu^{*}(K_{X})|_{F}\cdot\operatorname{Mov}|\sigma^{*}(K_{F_{0}})|)$ (which coincides with the definition in [7, §2.9]). Then by the first $3$ lines of [7, §2.15], one has $\xi\geq\frac{2}{3}$. By Lemma 3.3, $2\mu^{*}(K_{X})|_{F}-\sigma^{*}(K_{F_{0}})$ is effective. Then
$K_{X}^{3}={}$ $\mu^{*}(K_{X})^{3}\geq(\mu^{*}(K_{X})^{2}\cdot F)=(\mu^{*}(K_{X})|_{F})^{2}$
[…]
Then by assumption, all the above inequalities are equalities. ∎
Now we will analyze the geometry of $X$ in more details.
Lemma 4.2.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=2$ and $K_{X}^{3}=\frac{1}{3}$. Keep the notation in Setting 3.2. Then there exists a unique $\mu$-exceptional prime divisor $E_{0}$ such that $(\sigma^{*}(K_{F_{0}})\cdot E_{0}|_{F})>0$. Moreover,
  1. (1)
    $(\sigma^{*}(K_{F_{0}})\cdot E_{0}|_{F})=1$;
  2. (2)
    $\operatorname{coeff}_{E_{0}}E_{\mu}=\frac{1}{3}$, $\operatorname{coeff}_{E_{0}}Z=\frac{2}{3}$;
  3. (3)
    $\mu(E_{0})$ is a point.
Proof.
By the adjunction formula, we have
$K_{F}=\mu^{*}(K_{X})|_{F}+E_{\mu}|_{F}.$
By Lemma 4.1 and the equality $(K_{F}\cdot\sigma^{*}(K_{F_{0}}))=K_{F_{0}}^{2}=1$, we have
$(E_{\mu}|_{F}\cdot\sigma^{*}(K_{F_{0}}))=\frac{1}{3}.$ (4.1)
Thus there exists a $\mu$-exceptional prime divisor $E_{0}\subseteq\operatorname{Supp}E_{\mu}$ such that $(E_{0}|_{F}\cdot\sigma^{*}(K_{F_{0}}))>0,$ which means that $(E_{0}|_{F}\cdot\sigma^{*}(K_{F_{0}}))\geq 1.$ Hence
$((K_{X^{\prime}}-E_{0})|_{F}\cdot\sigma^{*}(K_{F_{0}}))=(K_{F}\cdot\sigma^{*}(K_{F_{0}}))-(E_{0}|_{F}\cdot\sigma^{*}(K_{F_{0}}))\leq 0.$ (4.2)
On the other hand,
$K_{X^{\prime}}=F+Z+E_{\mu}$
is an effective integral divisor whose support contains all $\mu$-exceptional divisors. So $K_{X^{\prime}}-E_{0}$ is an effective integral divisor. As $\sigma^{*}(K_{F_{0}})$ is nef, we conclude from (4.2) that
$((K_{X^{\prime}}-E_{0})|_{F}\cdot\sigma^{*}(K_{F_{0}}))=0;$ (4.3)
$(E_{0}|_{F}\cdot\sigma^{*}(K_{F_{0}}))=1.$ (4.4)
In particular, (4.3) implies that $\operatorname{coeff}_{E_{0}}(Z+E_{\mu})=1$ and $E_{0}$ is the unique $\mu$-exceptional prime divisor such that $(E_{0}|_{F}\cdot\sigma^{*}(K_{F_{0}}))>0$. Furthermore, (4.1) and (4.4) imply that $\operatorname{coeff}_{E_{0}}E_{\mu}=\frac{1}{3}$ and hence $\operatorname{coeff}_{E_{0}}Z=\frac{2}{3}$.
Finally, since $X$ has only isolated singularities and $\operatorname{coeff}_{E_{0}}E_{\mu}\not\in\mathbb{Z}$, we deduce that $\mu(E_{0})$ is a point. ∎
Lemma 4.3.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=2$ and $K_{X}^{3}=\frac{1}{3}$. Keep the notation in Setting 3.2 and Lemma 4.2. Then
  1. (1)
    $E_{0}|_{F}=C_{0}$ is an integral $(-3)$-curve;
  2. (2)
    $G|_{F}=\frac{1}{3}C_{0}$;
  3. (3)
    $Z_{X}=0$.
Proof.
By Lemma 4.2(1), there is a unique integral curve $C_{0}$ on $F$ with $\operatorname{coeff}_{C_{0}}E_{0}|_{F}=1$ such that $(\sigma^{*}(K_{F_{0}})\cdot C_{0})=1$. In particular, this implies that $C_{0}$ is not contracted by $\sigma$ and hence
$(K_{F}\cdot C_{0})\geq(\sigma^{*}(K_{F_{0}})\cdot C_{0})=1.$ (4.5)
Note that in Setting 3.2,
$Z-E_{\mu}=\mu^{*}(2K_{X})-K_{X^{\prime}}-F=\mu^{*}(Z_{X})+G$ (4.6)
is an effective $\mathbb{Q}$-divisor. Hence
$(Z-E_{\mu})|_{F}=\mu^{*}(2K_{X})|_{F}-K_{F}$ (4.7)
is an effective $\mathbb{Q}$-divisor with
$\operatorname{coeff}_{C_{0}}(Z-E_{\mu})|_{F}=\operatorname{coeff}_{E_{0}}(Z-E_{\mu})=\frac{1}{3}$
by Lemma 4.2(2). Hence by (4.7),
$\frac{1}{3}C_{0}^{2}\leq\big{(}(Z-E_{\mu})|_{F}\cdot C_{0}\big{)}=-(K_{F}\cdot C_{0}).$ (4.8)
Here we used the fact that $C_{0}$ is contracted by $\mu$ by Lemma 4.2(3). Combining (4.5) and (4.8), we have
$(K_{F}\cdot C_{0})+C_{0}^{2}\leq-2(K_{F}\cdot C_{0})\leq-2,$
According to the genus formula, we deduce that $(K_{F}\cdot C_{0})+C_{0}^{2}=-2$ and $C_{0}^{2}=-3$, which means that $C_{0}$ is a $(-3)$-curve.
Since $\sigma^{*}(K_{F_{0}})$ is nef and
$\left(\sigma^{*}(K_{F_{0}})+\frac{1}{3}C_{0}\right)\cdot C_{0}=1-1=0,$
we know that $\sigma^{*}(K_{F_{0}})+\frac{1}{3}C_{0}$ is nef. Moreover, $(\sigma^{*}(K_{F_{0}})+\frac{1}{3}C_{0})^{2}=\frac{4}{3}.$ On the other hand, from (4.7) we have
$\mu^{*}(2K_{X})|_{F}={}$ $K_{F}+(Z-E_{\mu})|_{F}\geq K_{F}+\frac{1}{3}E_{0}|_{F}$
[…]
Here $(\mu^{*}(2K_{X})|_{F})^{2}=\frac{4}{3}$ by Lemma 4.1. Hence by Lemma 2.2(1), all inequalities in (4.9) become equalities. In particular, $(Z-E_{\mu})|_{F}=\frac{1}{3}E_{0}|_{F}=\frac{1}{3}C_{0}$. Moreover, $(\mu^{*}(Z_{X})+G)|_{F}=\frac{1}{3}C_{0}$ by (4.6), so $\mu^{*}(Z_{X})|_{F}$ and $G|_{F}$ are multiples of $C_{0}$. Since $C_{0}$ is contracted by $\mu$ by Lemma 4.2(3), we have $\mu^{*}(Z_{X})|_{F}=0$ and $G|_{F}=\frac{1}{3}C_{0}$. The last assertion follows from Lemma 3.5. ∎
Lemma 4.4.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=2$ and $K_{X}^{3}=\frac{1}{3}$. Suppose that $P$ is a non-Gorenstein singularity of $X$. Then $r_{P}\in\{2,3\}$ and there is exactly one such $P$ with $r_{P}=3$. Here $r_{P}$ is the Cartier index of $K_{X}$ at $P$.
Proof.
By Lemma 4.3, all conditions of Lemma 3.4 are satisfied.
We claim that $F_{X}$ has exactly one non-Gorenstein singularity which is of type $\frac{1}{3}(1,1)$. Indeed, by Lemmas 4.3 and 3.4,
$K_{F}+\frac{1}{3}C_{0}=(\mu|_{F})^{*}(K_{F_{X}}).$
So outside the point $P_{0}=\mu|_{F}(C_{0})$, $F_{X}$ has at worst Du Val singularities which are clearly Gorenstein. The fiber of $\mu|_{F}$ over $P_{0}$ has simple normal crossing support as $F_{X}$ has only rational singularities, so the fiber of $\mu|_{F}$ over $P_{0}$ is exactly $C_{0}$, otherwise there exists an exceptional curve $C^{\prime}$ connecting $C_{0}$ with
$(K_{F}\cdot C^{\prime})=\frac{1}{3}(-C_{0}\cdot C^{\prime})=-\frac{1}{3},$
which is absurd. Since $C_{0}$ is a $(-3)$-curve, it is clear that $C_{0}$ is contracted to a cyclic quotient singularity of type $\frac{1}{3}(1,1)$.
Now to conclude the lemma, by Lemma 3.4(2), we only need to consider the case that $P=P_{0}$ is the unique non-Gorenstein singularity of $F_{X}$. Then $r_{P}=3$ as $3K_{X}|_{F_{X}}$ is Cartier at $P$. ∎
4.2. Geometry of pluricanonical systems
In this subsection, we study the pluricanonical maps of $X$.
Let $\pi:W\to X$ be as in Setting 2.1 and keep the same notation. We may further assume that $\pi$ factorizes through the modification $\mu:X^{\prime}\to X$ in §4.1. In this case, $S$ is a $(1,2)$-surface. Denote by $\sigma_{S}:S\to S_{0}=F_{0}$ the morphism to its minimal model. Let $C$ be a general member of $\operatorname{Mov}|K_{S}|=\operatorname{Mov}|\sigma_{S}^{*}K_{S_{0}}|$. Then $C$ is a smooth projective curve of genus $2$. Here for a good reference on geometry of $(1,2)$-surfaces, we refer to [16, §2].
Recall that
$(\pi^{*}(K_{X})|_{S})^{2}{}$ $=\frac{1}{3},$ (4.10)
$(\pi^{*}(K_{X})|_{S}\cdot C){}$ $=\frac{2}{3}$ (4.11)
by Lemma 4.1 and the projection formula.
Lemma 4.5.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=2$ and $K_{X}^{3}=\frac{1}{3}$. Then $P_{2}(X)\leq 4$.
Proof.
Keep the notation in Setting 2.1.
Step 1. We claim that $u_{2,0}\leq 2$.
Suppose that $u_{2,0}\geq 3$. If $|M_{2,0}||_{S}$ and $|C|$ are composed with the same pencil, then $M_{2,0}|_{S}\geq 2C$ since $u_{2,0}\geq 3$. Thus $2\pi^{*}(K_{X})|_{S}\geq 2C$, which implies that
$K_{X}^{3}\geq(\pi^{*}(K_{X})|_{S})^{2}\geq(\pi^{*}(K_{X})|_{S}\cdot C)=\frac{2}{3},$
which is a contradiction. On the other hand, if $|M_{2,0}||_{S}$ and $|C|$ are not composed with the same pencil, then $(M_{2,0}|_{S}\cdot C)\geq 2$ as $|M_{2,0}|$ induces a nontrivial movable linear system on $C$. Hence
$(\pi^{*}(K_{X})|_{S}\cdot C)\geq\frac{1}{2}(M_{2,0}|_{S}\cdot C)\geq 1,$
which contradicts (4.11). We conclude that $u_{2,0}\leq 2$.
Step 2. We claim that $u_{2,-1}\leq 1$.
Suppose that $u_{2,-1}\geq 2$. If $|M_{2,-1}||_{S}$ and $|C|$ are composed with the same pencil, then $M_{2,-1}|_{S}\geq C$ and hence
$2K_{X}^{3}$ $\geq(\pi^{*}(K_{X})^{2}\cdot S)+(\pi^{*}(K_{X})^{2}\cdot M_{2,-1})$
$\geq(\pi^{*}(K_{X})^{2}\cdot S)+(\pi^{*}(K_{X})|_{S}\cdot M_{2,-1}|_{S})$
$\geq(\pi^{*}(K_{X})^{2}\cdot S)+(\pi^{*}(K_{X})|_{S}\cdot C)$
$\geq 1,$
which is a contradiction. On the other hand, if $|M_{2,-1}||_{S}$ and $|C|$ are not composed with the same pencil, then $(M_{2,-1}|_{S}\cdot C)\geq 2$. Hence
$(\pi^{*}(K_{X})|_{S}\cdot C)\geq\frac{1}{2}(M_{2,-1}|_{S}\cdot C)\geq 1,$
which contradicts (4.11). Thus $u_{2,-1}\leq 1$.
Step 3. We claim that $h^{0}(W,M_{2,-2})\leq 1$.
Otherwise, $|M_{2,-2}|$ is a nontrivial movable linear system. Since $K_{X}$ is nef and big, we have $(\pi^{*}(K_{X})^{2}\cdot M_{2,-2})>0$. Recall that $2\pi^{*}(K_{X})-2S\geq M_{2,-2}$ by definition. Then by (4.10),
$K_{X}^{3}\geq(\pi^{*}(K_{X})|_{S})^{2}+\frac{1}{2}((\pi^{*}(K_{X}))^{2}\cdot M_{2,-2})>\frac{1}{3},$
which is a contradiction. Thus $h^{0}(W,M_{2,-2})\leq 1$.
Step 4. The conclusion that $P_{2}(X)\leq 4$.
This follows from
$P_{2}(X)=h^{0}(W,\lfloor{2\pi^{*}(K_{X})}\rfloor)=u_{2,0}+u_{2,-1}+h^{0}(W,M_{2,-2})$
and previous steps. ∎
Lemma 4.6.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=2$ and $K_{X}^{3}=\frac{1}{3}$. Then
  1. (1)
    $P_{3}(X)\geq P_{2}(X)+3$;
  2. (2)
    $|3K_{X}|$ defines a generically finite map;
  3. (3)
    $P_{4}(X)\geq P_{3}(X)+4$.
Proof.
Keep the notation in Setting 2.1.
By the Kawamata–Viehweg vanishing theorem, the natural restriction map
$H^{0}(W,K_{W}+\lceil{\pi^{*}(K_{X})}\rceil+S)\to H^{0}(S,K_{S}+\lceil{\pi^{*}(K_{X})|_{S}}\rceil)$
is surjective. Here in the last term, the restriction on $S$ commutes with the roundup as $S$ is general. So
$|3K_{W}||_{S}\succcurlyeq|K_{W}+\lceil{\pi^{*}(K_{X})}\rceil+S||_{S}\succcurlyeq|K_{S}+\lceil{\pi^{*}(K_{X})|_{S}}\rceil|.$ (4.12)
Note that $2\pi^{*}(K_{X})|_{S}\geq\sigma_{S}^{*}(K_{S_{0}})$ by Lemma 3.3 as $\pi$ factorizes through $\mu$. By [2, Theorem 3.2], we have $h^{0}(S,K_{S}+\lceil{\pi^{*}(K_{X})|_{S}}\rceil)\geq 3$. Then $u_{3,0}\geq 3$, which implies that
$P_{3}(X)=h^{0}(W,\lfloor{3\pi^{*}(K_{X})}\rfloor-S)+u_{3,0}\geq P_{2}(X)+3.$
Assertion (1) is proved.
Next, we claim that $|K_{S}+\lceil{\pi^{*}(K_{X})|_{S}}\rceil|$ defines a generically finite map on $S$. Otherwise, it is composed with the same pencil with $|C|$ and thus $K_{S}+\lceil{\pi^{*}(K_{X})|_{S}}\rceil\geq 2C$ since $h^{0}(S,K_{S}+\lceil{\pi^{*}(K_{X})|_{S}}\rceil)\geq 3$. By comparing the movable parts of both sides of (4.12), we know that $3\pi^{*}(K_{X})|_{S}\geq 2C$. Thus by (4.11),
$K_{X}^{3}\geq(\pi^{*}(K_{X})|_{S})^{2}\geq\frac{2}{3}(\pi^{*}(K_{X})|_{S}\cdot C)=\frac{4}{9},$
which is a contradiction. So $|K_{S}+\lceil{\pi^{*}(K_{X})|_{S}}\rceil|$ defines a generically finite map on $S$. By (4.12), $|3K_{X}|$ defines a generically finite map. Assertion (2) is proved.
By [10, Theorem 2.4(1)], the natural restriction map
$H^{0}(W,2(K_{W}+S))\to H^{0}(S,2K_{S})$
is surjective. Note that $P_{2}(S)=4$ and $|4K_{W}|\succcurlyeq|2(K_{W}+S)|$. We deduce that $u_{4,0}\geq 4$, which implies that $P_{4}(X)\geq P_{3}(X)+4$. ∎
Proposition 4.7.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=2$ and $K_{X}^{3}=\frac{1}{3}$. Then
  1. (1)
    $\chi(\mathcal{O}_{X})=-1$, $P_{2}(X)=4$;
  2. (2)
    the Reid basket of $X$ is $\{2\times(1,2),(1,3)\}$;
  3. (3)
    the Hilbert series of $X$ is given by
    $\sum_{k\geq 0}h^{0}(X,kK_{X})q^{k}=\frac{1-q^{16}}{(1-q)^{2}(1-q^{2})(1-q^{3})(1-q^{8})}.$
All the above data are the same as those of a general hypersurface of degree $16$ in $\mathbb{P}(1,1,2,3,8)$.
Proof.
By Lemma 4.4 and [30, (6.1), (6.4)], the basket $B_{X}$ of $X$ is $\{a\times(1,2),b\times(1,3)\}$, for some integers $a\geq 0$ and $b\geq 1$. By Reid’s Riemann–Roch formula (2.1) and the Kawamata–Viehweg vanishing theorem,
$P_{2}(X){}$ $=\frac{1}{6}-3\chi(\mathcal{O}_{X})+\frac{a}{4}+\frac{b}{3};$
$P_{3}(X){}$ $=\frac{5}{6}-5\chi(\mathcal{O}_{X})+\frac{a}{4}+\frac{2b}{3};$
$P_{4}(X){}$ $=\frac{7}{3}-7\chi(\mathcal{O}_{X})+\frac{a}{2}+\frac{2b}{3}.$
Then it is clear that
$2P_{4}(X)-P_{3}(X)-3P_{2}(X)=\frac{10}{3}-\frac{b}{3}\leq 3$
On the other hand, by Lemmas 4.6 and 4.5,
$2P_{4}(X)-P_{3}(X)-3P_{2}(X)=P_{4}(X)-P_{3}(X)+P_{4}(X)-P_{2}(X)-2P_{2}(X)\geq 3$
Then all the above inequalities become equalities, which means that $b=1$, $P_{2}(X)=4$, $P_{3}(X)=7$, $P_{4}(X)=11$. Then we can solve the equations and get $\chi(\mathcal{O}_{X})=-1$ and $a=2$. Assertions (1) and (2) are proved.
Note that Reid’s Riemann–Roch formula only depends on $K_{X}^{3}$, $\chi(\mathcal{O}_{X})$ and $B_{X}$. These data are the same as those of a general hypersurface of degree $16$ in $\mathbb{P}(1,1,2,3,8)$ (see [18, Table 3, No. 12]). So the Hilbert series of them also coincide and are certainly computable (see [13, Theorem 3.4.4]). ∎
Proposition 4.8.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=2$ and $K_{X}^{3}=\frac{1}{3}$. Then
  1. (1)
    $|2K_{X}|$ is not composed with a pencil;
  2. (2)
    $|3K_{X}|$ defines a generically finite map;
  3. (3)
    $|8K_{X}|$ defines a birational map.
Proof.
If $|2K_{X}|$ is composed with a pencil, then it is the same pencil as $|K_{X}|$ and $2\pi^{*}(K_{X})\geq 3S$ since $P_{2}(X)=4$. Thus by (4.10),
$K_{X}^{3}\geq\frac{3}{2}(\pi^{*}(K_{X})|_{S})^{2}\geq\frac{1}{2},$
which is a contradiction. This proves Assertion (1).
Assertion (2) follows from Lemma 4.6. Assertion (3) follows from [7, Theorem 1.4]. ∎
5. Birational geometry of $3$-folds with $p_{g}=3$ and $\text{\rm Vol}=1$
In this section, we study the birational geometry of a minimal $3$-fold $X$ of general type with $p_{g}(X)=3$ and $K_{X}^{3}=1$.
5.1. Singularities
In this subsection, we study the Gorenstein index of $X$.
Take a birational modification $\mu:X^{\prime}\to X$ as in Setting 3.2 and keep the same notation. By [7, Theorem 1.5(4)], the canonical image of $X$ is $\mathbb{P}^{2}$. By [8, Lemma 5.2], $F$ is a $(4,4)$-surface.
The geometry of $X$ is simpler than that in §4.
Lemma 5.1.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=3$ and $K_{X}^{3}=1$. Keep the notation in Setting 3.2. Then
  1. (1)
    $\mu^{*}(K_{X})|_{F}\equiv\frac{1}{2}K_{F}\equiv\frac{1}{2}\sigma^{*}(K_{F_{0}})$; in particular, $F$ is minimal;
  2. (2)
    $G|_{F}=0$, $Z_{X}=0$.
Proof.
By (3.1), we have
$K_{F}+G|_{F}=\mu^{*}(K_{X}+F_{X})|_{F}.$
Since $G$ is effective, we deduce that
$2\mu^{*}(K_{X})|_{F}=K_{F}+G|_{F}+\mu^{*}(Z_{X})|_{F}\geq K_{F}\geq\sigma^{*}(K_{F_{0}}).$ (5.1)
Then $(\mu^{*}(K_{X})|_{F})^{2}\geq\frac{1}{4}K_{F_{0}}^{2}=1$. On the other hand, $(\mu^{*}(K_{X})|_{F})^{2}\leq K_{X}^{3}=1$. Hence $(\mu^{*}(K_{X})|_{F})^{2}=\frac{1}{4}K_{F_{0}}^{2}=1$. By Lemma 2.2(1), inequalities in (5.1) become equalities. In particular, Assertion (1) holds and $G|_{F}=\mu^{*}(Z_{X})|_{F}=0$. The last assertion follows from Lemma 3.5. ∎
Lemma 5.2.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=3$ and $K_{X}^{3}=1$. Suppose that $P$ is a non-Gorenstein singularity of $X$. Then $r_{P}=2$. Here $r_{P}$ is the Cartier index of $K_{X}$ at $P$.
Proof.
By Lemma 5.1, all conditions of Lemma 3.4 are satisfied. Moreover, $G|_{F}=0$ implies that $K_{F}=(\mu|_{F})^{*}(K_{F_{X}})$. So $F_{X}$ has at worst Du Val singularities which are clearly Gorenstein, which implies that $r_{P}=2$ by Lemma 3.4(2). ∎
5.2. Geometry of pluricanonical systems
Take a birational modification $\pi:W\to X$ as in Setting 2.1 and keep the same notation. We may further assume that $\pi$ factorizes through the modification $\mu:X^{\prime}\to X$ in §5.1. In this case, $S$ is a $(4,4)$-surface. By [7, §3.2], $|S|_{S}|$ is composed with a free rational pencil and $S|_{S}\sim C$, where $C$ is a general fiber of $\psi:W\to\Sigma^{\prime}$.
Lemma 5.3.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=3$ and $K_{X}^{3}=1$. Then $P_{2}(X)\leq 7$. Moreover, the equality holds only if $|2K_{X}|$ defines a generically finite map of degree $2$.
Proof.
Recall that $\Phi_{|2K_{X}|}$ denotes the rational map defined by $|2K_{X}|$. Denote by $\Sigma_{2}$ the image of $\Phi_{|2K_{X}|}$. Since the canonical image of $X$ is $\mathbb{P}^{2}$ by [7, Theorem 1.5(4)], it is clear that $\dim\Sigma_{2}\geq 2$.
If $\dim\Sigma_{2}=2$, then $|2K_{X}|$ induces a fibration $\psi_{2}:W\to\Sigma^{\prime}_{2}$ which factors through $\psi:W\to\Sigma^{\prime}$ and $\Sigma^{\prime}_{2}$ is birational to $\Sigma^{\prime}$. In particular, we can identify a general fiber of $\psi_{2}$ with $C$. So we may write ${M_{2}}|_{M_{2}}\equiv d_{2}C$, where $M_{2}$ is a general element in $\operatorname{Mov}|2\pi^{*}(K_{X})|$ and $d_{2}\geq P_{2}(X)-2$. Then
$4K_{X}^{3}\geq(\pi^{*}(K_{X})\cdot M_{2}\cdot M_{2})\geq d_{2}(\pi^{*}(K_{X})\cdot C)\geq d_{2}\geq P_{2}(X)-2,$
where the third inequality follows from [7, §3.2] as $(\pi^{*}(K_{X})\cdot C)=\xi\geq 1$. We deduce that $P_{2}(X)\leq 6$ when $\dim\Sigma_{2}=2$.
If $\dim\Sigma_{2}=3$, then by [7, Theorem 1.5(4)], $\Phi_{|4K_{X}|}$ is not birational. So $\deg(\Phi_{|2K_{X}|})\geq 2$. Then
$8K_{X}^{3}\geq\deg(\Phi_{|2K_{X}|})\deg(\Sigma_{2})\geq 2(P_{2}(X)-3),$
where $\deg(\Sigma_{2})\geq P_{2}(X)-3$. We conclude that $P_{2}(X)\leq 7$. Moreover, when $P_{2}(X)=7$, the above inequalities become equalities. In particular, $\deg(\Phi_{|2K_{X}|})=2$. ∎
Lemma 5.4.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=3$ and $K_{X}^{3}=1$. Then $P_{3}(X)\geq P_{2}(X)+6$.
Proof.
Kepp the notation in Setting 2.1. By the Kawamata–Viehweg vanishing theorem, we have
$|3K_{W}||_{S}\succcurlyeq|K_{W}+\lceil{\pi^{*}(K_{X})}\rceil+S||_{S}\succcurlyeq|K_{S}+\lceil{\pi^{*}(K_{X})|_{S}}\rceil|.$
Since $\pi^{*}(K_{X})|_{S}\geq S|_{S}=C$, we have
$|3K_{W}||_{S}\succcurlyeq|K_{S}+C|.$ (5.2)
By [15, Theorem 3.4], $h^{1}(\mathcal{O}_{S})=0$. By the exact sequence
$0\to\mathcal{O}_{S}(K_{S})\to\mathcal{O}_{S}(K_{S}+C)\to\mathcal{O}_{C}(K_{C})\to 0,$
we deduce that $h^{0}(S,K_{S}+C)=p_{g}(S)+g(C)\geq 6$. Thus (5.2) implies that
$P_{3}(X)\geq h^{0}(3K_{W}-S)+h^{0}(S,K_{S}+C)\geq P_{2}(X)+6.$
The proof is completed. ∎
Proposition 5.5.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=3$ and $K_{X}^{3}=1$. Then
  1. (1)
    $\chi(\mathcal{O}_{X})=-2$, $P_{2}(X)=7$;
  2. (2)
    the basket of $X$ is $B_{X}=\{2\times(1,2)\}$;
  3. (3)
    the Hilbert series of $X$ is given by
    $\sum_{k\geq 0}h^{0}(X,kK_{X})q^{k}=\frac{1-q^{12}}{(1-q)^{3}(1-q^{2})(1-q^{6})}.$
All the above data are the same as those of a general hypersurface of degree $12$ in $\mathbb{P}(1,1,1,2,6)$.
Proof.
By Lemma 5.2 and [30, (6.1), (6.4)], the basket $B_{X}$ of $X$ is $\{a\times(1,2)\}$ for some $a\geq 0$. By Reid’s Riemann–Roch formula (2.1) and the Kawamata–Viehweg vanishing theorem, we have
$P_{2}(X){}$ $=\frac{1}{2}-3\chi(\mathcal{O}_{X})+\frac{a}{4},$
$P_{3}(X){}$ $=\frac{5}{2}-5\chi(\mathcal{O}_{X})+\frac{a}{4}.$
Then by Lemmas 5.4 and 5.3,
$14\geq 2P_{2}(X)=3P_{3}(X)-3P_{2}(X)-5+\frac{a}{2}\geq 13+\frac{a}{2}.$
So $P_{2}(X)=7$ and $a\leq 2$. Then it is not hard to see that $a=2$ and $\chi(\mathcal{O}_{X})=-2$ from the equation of $P_{2}(X)$. Assertions (1) and (2) are proved.
Note that Reid’s Riemann–Roch formula only depends on $K_{X}^{3}$, $\chi(\mathcal{O}_{X})$ and $B_{X}$. These data are the same as those of a general hypersurface of degree $12$ in $\mathbb{P}(1,1,1,2,6)$ (see [18, Table 3, No. 7]). So the Hilbert series of them also coincide and can be computed by [13, Theorem 3.4.4]. ∎
Proposition 5.6.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=3$ and $K_{X}^{3}=1$. Then
  1. (1)
    $|K_{X}|$ is not composed with a pencil;
  2. (2)
    $|2K_{X}|$ defines a generically finite map of degree $2$;
  3. (3)
    $|6K_{X}|$ defines a birational map.
Proof.
Assertion (1) follows from [7, Theorem 1.5(4)]. Assertion (2) follows from Lemma 5.3 and $P_{2}(X)=7$ in Proposition 5.5. Assertion (3) follows from [6, Theorem 1.2]. ∎
6. Birational geometry of $3$-folds with $p_{g}=4$ and $\text{\rm Vol}=2$
In this section, we study the birational geometry of a minimal $3$-fold $X$ of general type with $p_{g}(X)=4$ and $K_{X}^{3}=2$.
Recall that $\Sigma$ denotes the canonical image of $X$ in $\mathbb{P}^{3}$. By [7, Theorem 1.5(5)], we have $\dim\Sigma\geq 2$.
When $\dim\Sigma=3$, the geometry of $X$ is well-understood by Kobayashi [22]. We summarize the result with more details as the following.
Proposition 6.1.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=4$ and $K_{X}^{3}=2$. Suppose that the canonical map of $X$ is generically finite. Then
  1. (1)
    $X$ is Gorenstein;
  2. (2)
    the canonical map is a generically finite morphism to $\mathbb{P}^{3}$ of degree $2$;
  3. (3)
    the canonical model of $X$ is a hypersurface of degree $10$ in $\mathbb{P}(1,1,1,1,5)$;
  4. (4)
    the Hilbert series of $X$ is given by
    $\sum_{k\geq 0}h^{0}(X,kK_{X})q^{k}=\frac{1-q^{10}}{(1-q)^{4}(1-q^{5})}.$
Proof.
Assertions (1) and (2) are due to [22, Proposition 2.5]. Denote by $Y$ the canonical model of $X$ and $\phi:Y\to\mathbb{P}^{3}$ the canonical map which is a morphism defined by $|K_{Y}|$. Then $\phi$ is finite as $K_{Y}$ is ample. Moreover, $\deg(\phi)=2$. Denote by $D\subseteq\mathbb{P}^{3}$ the branch locus, then $K_{Y}=\phi^{*}(K_{\mathbb{P}^{3}}+\frac{1}{2}D)$ by the ramification formula. By computing the self-intersection number, we get $D\in|\mathcal{O}_{\mathbb{P}^{3}}(10)|$. So $Y$ can be expressed as a hypersurface of degree $10$ in $\mathbb{P}(1,1,1,1,5)$ defined by $x_{4}^{2}+f_{0}(x_{0},x_{1},x_{2},x_{3})=0$, where $[x_{0}:\dots:x_{4}]$ are coordinates of $\mathbb{P}(1,1,1,1,5)$ and $f_{0}$ is the defining equation of $D\subseteq\mathbb{P}^{3}$. Finally, to compute the Hilbert series, we can use either [13, 3.4.2] or Reid’s Riemann–Roch formula (2.1). Here recall that $K_{X}^{3}=2$ and $B_{X}=\emptyset$, and $\chi({\mathcal{O}}_{X})=-3$ which can be computed from the double cover construction. ∎
Thus we only need to treat the case when $\dim\Sigma=2$. We recall the following properties from [7, §3.2].
Lemma 6.2.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=4$ and $K_{X}^{3}=2$. Suppose that the canonical map of $X$ is not generically finite. Let $\pi:W\to X$ be as in Setting 2.1 and keep the same notation. Denote by $C$ a general fiber of $\psi:W\to\Sigma^{\prime}$. Then
  1. (1)
    $(\pi^{*}(K_{X})\cdot C)=1$;
  2. (2)
    $C$ is a smooth projective curve of genus $2$;
  3. (3)
    $S|_{S}\sim 2C$ which is a rational pencil;
  4. (4)
    $\Sigma^{\prime}\simeq\Sigma\subseteq\mathbb{P}^{3}$ is a nondegenerate surface of degree $2$, in particular, $\Sigma$ is either $\mathbb{P}(1,1,2)$ or $\mathbb{P}^{1}\times\mathbb{P}^{1}$.
Proof.
Assertions (1)–(3) are directly from [7, §3.2]. For Assertion (4), just note that $S|_{S}\equiv\deg(s)\deg(\Sigma)C$ and $\deg(\Sigma)\geq 2$ as it is nondegenerate. So $\deg(\Sigma)=2$ and $\deg(s)=1$ by Assertion (3). By [28, §10], $\Sigma$ is either $\mathbb{P}(1,1,2)$ or $\mathbb{P}^{1}\times\mathbb{P}^{1}$. In particular, $\Sigma$ is normal and hence $s$ is an isomorphism. ∎
By adopting the idea in [17, Proof of Proposition 2.1], we have the following lemma.
Lemma 6.3.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=4$ and $K_{X}^{3}=2$. Suppose that the canonical map of $X$ is not generically finite. Then there exits a minimal $3$-fold $X_{1}$ birational to $X$, admitting a fibration $f_{1}:X_{1}\to\mathbb{P}^{1}$ whose general fibers are $(1,2)$-surfaces.
Proof.
By Lemma 6.2, $\Sigma$ is either $\mathbb{P}(1,1,2)$ or $\mathbb{P}^{1}\times\mathbb{P}^{1}$.
Case 1. $\Sigma\simeq\mathbb{P}(1,1,2)$.
In this case, there is a contraction $r:\mathbb{F}_{2}\to\mathbb{P}^{3}$ from the second Hirzebruch surface induced by the linear system $|s+2\ell|$ such that $\Sigma=r(\mathbb{F}_{2})$. Here $\ell$ is the ruling of the natural fibration $p:\mathbb{F}_{2}\to\mathbb{P}^{1}$ and $s$ is the unique negative section.
Replacing $W$ by its birational model, we may assume that there is a morphism $\varphi:W\to\mathbb{F}_{2}$ such that $\psi=r\circ\varphi$. Thus we obtain a fibration $f_{W}:=p\circ\varphi:W\to\mathbb{P}^{1}$ with a general fiber $F_{W}=\varphi^{*}(\ell)$. Let $\zeta:W\dashrightarrow X_{1}$ be the contraction to a relative minimal model over $\mathbb{P}^{1}$, where we have a relatively minimal fibration $f_{1}:X_{1}\to\mathbb{P}^{1}$ with a general fiber denoted by $F_{1}$. Up to a further birational modification of $W$, we may assume that $\zeta$ is a morphism. Thus $\sigma_{1}:=\zeta|_{F_{W}}:F_{W}\to F_{1}$ is just the contraction to its minimal model. We claim that $F_{1}$ is a $(1,2)$-surface and $X_{1}$ is minimal.
Note that by definition,
$\pi^{*}(K_{X})\geq\psi^{*}\mathcal{O}_{\Sigma}(1)\geq\varphi^{*}(2\ell)=2F_{W}.$
By [4, Corollary 2.3], $\pi^{*}(K_{X})|_{F_{W}}-\frac{2}{3}\sigma_{1}^{*}(K_{F_{1}})$ is $\mathbb{Q}$-effective. Note that $S|_{F_{W}}\sim C$. By Lemma 6.2(1), we have
$(\sigma_{1}^{*}(K_{F_{1}})\cdot C)\leq\frac{3}{2}(\pi^{*}(K_{X})|_{F_{W}}\cdot C)=\frac{3}{2}.$
Then $(\sigma_{1}^{*}(K_{F_{1}})\cdot C)=1$. Since $C$ is a moving curve on $F_{W}$, we conclude that $F_{1}$ is a $(1,2)$-surface by [3, Lemma 2.4]. Note that $K_{F_{W}}\geq\pi^{*}(K_{X})|_{F_{W}}$ and by Lemma 6.2(1),
$(\pi^{*}(K_{X})|_{F_{W}})^{2}\geq(\pi^{*}(K_{X})|_{F_{W}}\cdot S|_{F_{W}})=(\pi^{*}(K_{X})|_{F_{W}}\cdot C)=1=(\sigma_{1}^{*}(K_{F_{1}}))^{2}.$
We deduce that $\pi^{*}(K_{X})|_{F_{W}}=\sigma_{1}^{*}(K_{F_{1}})$ by considering the Zariski decomposition of $K_{F_{W}}$ and Lemma 2.2(1). Thus $X_{1}$ is minimal by [4, Lemma 3.2]. We finish the proof in this case.
Case 2. $\Sigma\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}$.
Denote by $p_{1}$ and $p_{2}$ the two natural projections of $\mathbb{P}^{1}\times\mathbb{P}^{1}$. For $i=1,2$, write $f_{W,i}=p_{i}\circ\gamma:W\to\mathbb{P}^{1}$. Then $f_{W,i}$ is a fibration for each $i$. Let $T_{i}$ be a general fiber of $f_{W,i}$. Then
$\pi^{*}(K_{X})\geq\psi^{*}(\mathcal{O}_{\Sigma}(1))=T_{1}+T_{2}.$ (6.1)
Since $K_{X}^{3}=2$, we may assume that $(\pi^{*}(K_{X})|_{T_{1}})^{2}\leq 1$. Same as in Case 1, we get a contraction $\zeta:W\to X_{1}$ to a relative minimal model of $f_{W,1}$, where we denote by $f_{1}:X_{1}\to\mathbb{P}^{1}$ the relatively minimal fibration and by $F_{1}$ its general fiber. Thus $\sigma_{1}:=\zeta|_{T_{1}}:T_{1}\to F_{1}$ is just the contraction to its minimal model. We claim that $F_{1}$ is a $(1,2)$-surface and $X_{1}$ is minimal.
For sufficiently large $m$, we have
$|2mK_{W}||_{T_{1}}\succcurlyeq|m(K_{W}+T_{1}+T_{2})||_{T_{1}}\succcurlyeq|m(K_{T_{1}}+{T_{2}}|_{T_{1}})|,$
where the first inequality is by (6.1) and the second is by [10, Theorem 2.4(2)]. Hence by considering the movable parts of both sides, we deduce that $\pi^{*}(2K_{X})|_{T_{1}}-\sigma_{1}^{*}(K_{F_{1}})-{T_{2}}|_{T_{1}}$ is $\mathbb{Q}$-effective. Note that $T_{2}|_{T_{1}}=C$. Combining with Lemma 6.2(1) and taking the intersection with $\pi^{*}(K_{X})|_{T_{1}}$, we get
$(\pi^{*}(K_{X})|_{T_{1}}\cdot\sigma_{1}^{*}(K_{F_{1}}))\leq 2(\pi^{*}(K_{X})|_{T_{1}})^{2}-(\pi^{*}(K_{X})|_{T_{1}}\cdot C)\leq 1.$
Since $\pi^{*}(K_{X})|_{T_{1}}\geq C$ by (6.1), we have $(C\cdot\sigma_{1}^{*}(K_{F_{1}}))\leq 1$. By [3, Lemma 2.4], $F_{1}$ is a $(1,2)$-surface. Furthermore, by Lemma 6.2(1),
$11\geq(\pi^{*}(K_{X})|_{T_{1}})^{2}\geq(\pi^{*}(K_{X})|_{T_{1}}\cdot C)=$ 1 .
Thus $(\pi^{*}(K_{X})|_{T_{1}})^{2}=1$. By considering the Zariski decomposition of $K_{T_{1}}$ and Lemma 2.2(1), we have $\pi^{*}(K_{X})|_{T_{1}}=\sigma_{1}^{*}(K_{F_{1}})$. By [4, Lemma 3.2], $X_{1}$ is minimal. The proof is completed. ∎
Proposition 6.4.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)=4$ and $K_{X}^{3}=2$. Suppose that the canonical map of $X$ is not generically finite. Then
  1. (1)
    $h^{1}(\mathcal{O}_{X})=h^{2}(\mathcal{O}_{X})=0$; in particular, $\chi(\mathcal{O}_{X})=-3$;
  2. (2)
    $X$ is Gorenstein;
  3. (3)
    the Hilbert series of $X$ is given by
    $\sum_{k\geq 0}h^{0}(X,kK_{X})q^{k}=\frac{1-q^{10}}{(1-q)^{4}(1-q^{5})};$
  4. (4)
    $|2K_{X}|$ defines a generically finite map of degree $2$;
  5. (5)
    $|5K_{X}|$ defines a birational map.
Proof.
By Lemma 6.3, after replacing by a birational model, we may assume that $X$ admits a fibration $f:X\to\mathbb{P}^{1}$ with a general fiber $F$ a $(1,2)$-surface. Now we are in the setting of [17, §3]. Assertion (1) follows from [17, Lemma 3.4].
Recall Reid’s Riemann–Roch formula (2.1) for $P_{2}(X)$:
$P_{2}(X)=\frac{1}{2}K_{X}^{3}-3\chi(\mathcal{O}_{X})+l_{2}(X).$
Here $l_{2}(X)\geq 0$ and $l_{2}(X)=0$ if and only if $X$ is Gorenstein. So $P_{2}(X)\geq 10$. On the other hand, since $K_{X}^{3}<\frac{4}{3}p_{g}(X)-\frac{17}{6}$, by [17, Proposition 3.7], we have
$P_{2}(X)\leq\lfloor{2K_{X}^{3}}\rfloor+\lfloor{2K_{X}^{3}-\frac{5(p_{g}(X)-1)}{3}}\rfloor+7=10.$
So $P_{2}(X)=10$. Thus $l_{2}(X)=0$ and $X$ is Gorenstein. Assertion (2) is proved.
For Assertion (3), note that Reid’s Riemann–Roch formula only depends on $K_{X}^{3}$, $\chi(\mathcal{O}_{X})$ and $B_{X}$. These data are the same as those of a general hypersurface of degree $10$ in $\mathbb{P}(1,1,1,1,5)$ (see [18, Table 3, No. 5]). So the Hilbert series of them also coincide and can be computed by [13, Theorem 3.4.4].
For Assertion (4), note that the inequality in [17, Proposition 3.7(1)] is an equality now. The inequality in Step 1 of [17, Proof of Proposition 3.7] is also an equality. In particular, $|2K_{W}||_{F}$ defines a generically finite map of degree at most $2$. Thus $|2K_{X}|$ defines a generically finite map of degree at most $2$. However, $|2K_{X}|$ does not define a birational map as $|2K_{F}|$ does not define a birational map on a $(1,2)$-surface $F$ by [16, Lemma 2.1]. So $|2K_{X}|$ defines a generically finite map of degree $2$.
Assertion (5) follows from [6, Theorem 1.2]. ∎
7. Characterization of weighted hypersurfaces
The following theorem is a refinement of [20]. It provides a criterion for embedding a polarized variety into a weighted projective space by numerical and geometric properties of this variety.
Theorem 7.1.
Let $Y$ be a normal projective $3$-fold and let $H$ be a $\mathbb{Q}$-Cartier Weil ample divisor on $Y$. Suppose that there exists a well-formed quasismooth weighted hypersurface
$X_{d}\subseteq\mathbb{P}(1,a_{1},a_{2},a_{3},a_{4})$
with $1\leq a_{1}\leq a_{2}\leq a_{3}<a_{4}$ and $d=2a_{4}$ such that
  1. (1)
    $H^{3}=(\mathcal{O}_{X_{d}}(1))^{3}=\frac{2}{a_{1}a_{2}a_{3}}$ and
  2. (2)
    $h^{0}(Y,kH)=h^{0}(X_{d},\mathcal{O}_{X_{d}}(k))$ for any nonnegative integer $k$.
Suppose further that
  1. (3)
    $|a_{2}H|$ is not composed with a pencil;
  2. (4)
    $|a_{3}H|$ defines a generically finite map, while either the degree of this map is greater than $1$, or $Y$ is not rational;
  3. (5)
    $|a_{4}H|$ defines a birational map.
Then $Y$ is a hypersurface in $\mathbb{P}(1,a_{1},a_{2},a_{3},a_{4})$ defined by a weighted homogeneous polynomial $f$ of degree $d$, where
$f(x_{0},\dots,x_{4})=x_{4}^{2}+f_{0}(x_{0},x_{1},x_{2},x_{3})$
in suitable homogeneous coordinates $[x_{0}:\dots:x_{4}]$ of $\mathbb{P}(1,a_{1},a_{2},a_{3},a_{4})$; in other words, there is an isomorphism between graded $\mathbb{C}$-algebras
$\bigoplus_{k\geq 0}H^{0}(Y,kH)\simeq\mathbb{C}[x_{0},\dots,x_{4}]/(f),$
where $\mathbb{C}[x_{0},\dots,x_{4}]$ is viewed as a weighted polynomial ring with
$\operatorname{wt}(x_{0},\dots,x_{4})=(1,a_{1},a_{2},a_{3},a_{4}).$
Remark 7.2.
In Theorem 7.1, by [13, 3.4.2], we have the following formula:
$\sum_{k\geq 0}h^{0}(X_{d},\mathcal{O}_{X_{d}}(k))q^{k}=\frac{1-q^{d}}{(1-q)(1-q^{a_{1}})(1-q^{a_{2}})(1-q^{a_{3}})(1-q^{a_{4}})}.$
We will always use this formula to compute $h^{0}(Y,kH)$. In practice, we do not need to calculate the precise value of $h^{0}(Y,kH)$, we will often identify the value with the cardinality of a certain set in a combinatorial way instead.
Lemma 7.3.
Keep the notation in Theorem 7.1. For $0\leq i\leq 4$, take distinct general elements $g_{i}\in H^{0}(Y,a_{i}H)\setminus\{0\}$ (where $a_{0}=1$). For a positive integer $k$, denote
$S_{k}=\{g_{0}^{s_{0}}g_{1}^{s_{1}}g_{2}^{s_{2}}g_{3}^{s_{3}}\mid s_{0},\dots,s_{3}\in\mathbb{Z}_{\geq 0},s_{0}+a_{1}s_{1}+a_{2}s_{2}+a_{3}s_{3}=k\}.$
Then
  1. (1)
    the set $\{g_{0},g_{1},g_{2},g_{3}\}$ is algebraically independent in the graded algebra $R(Y,H)=\bigoplus_{k\geq 0}H^{0}(Y,kH);$
  2. (2)
    $H^{0}(Y,a_{4}H)$ is spanned by a basis $S_{a_{4}}\cup\{g_{4}\}.$
Here we remark that $R(Y,H)$ is viewed as a subalgebra of $\mathbb{C}(Y)[u]$, where the variable $u$ is from the grading of $R(Y,H)$. So a rational function $g\in H^{0}(Y,kH)\subseteq\mathbb{C}(Y)$ is identified with $gu^{k}\in R(Y,H)$. But to simplify the notation, we will just write $g\in R(Y,H)$ as long as the grading of $g$ is clear in the context.
Proof.
First we show the following claim.
Claim 7.4.
The set $\{g_{0},g_{1},g_{2}\}$ is algebraically independent in $R(Y,H)$.
Proof.
As $h^{0}(Y,a_{1}H)\geq 2$, $\{g_{0},g_{1}\}$ is algebraically independent. By assumption, $|a_{2}H|$ is not composed with a pencil. So by the generality of $g_{2}$, the transcendental degree of $\mathbb{C}(g_{0},g_{1},g_{2})$ as a subfield of the fractional field of $R(Y,H)$ is greater than $2$, which implies that $\{g_{0},g_{1},g_{2}\}$ is algebraically independent. ∎
By Claim 7.4,
$S_{k}^{\prime}=\{g_{0}^{s_{0}}g_{1}^{s_{1}}g_{2}^{s_{2}}\mid s_{0},s_{1},s_{2}\in\mathbb{Z}_{\geq 0},s_{0}+a_{1}s_{1}+a_{2}s_{2}=k\}$
is linearly independent in $H^{0}(Y,kH)$ for any positive integer $k$. On the other hand,
$h^{0}(Y,a_{3}H)=h^{0}(X_{d},\mathcal{O}_{X_{d}}(a_{3}))=|S_{a_{3}}|=|S^{\prime}_{a_{3}}|+1.$
Then $S_{a_{3}}=S^{\prime}_{a_{3}}\cup\{g_{3}\}$ is a basis of $H^{0}(Y,a_{3}H)$ by the generality of $g_{3}$. By assumption, $|a_{3}H|$ defines a generically finite map, so the transcendental degree of $\mathbb{C}(g_{0},g_{1},g_{2},g_{3})$ as a subfield of the fractional field of $R(Y,H)$ is greater than $3$, which implies that $\{g_{0},g_{1},g_{2},g_{3}\}$ is algebraically independent. This proves Assertion (1).
In particular, $S_{k}$ is linearly independent in $H^{0}(Y,kH)$ for any positive integer $k$. So Assertion (2) follows from the computation that
$h^{0}(Y,a_{4}H)=h^{0}(X_{d},\mathcal{O}_{X_{d}}(a_{4}))=|S_{a_{4}}|+1$
and the generality of $g_{4}$. ∎
Proof of Theorem 7.1.
Keep the notation in Lemma 7.3. We can define $3$ rational maps:
$\Psi_{a_{2}}:{}$ $Y\dashrightarrow\mathbb{P}(1,a_{1},a_{2});$
$P\mapsto[g_{0}(P):g_{1}(P):g_{2}(P)];$
$\Psi_{a_{3}}:{}$ $Y\dashrightarrow\mathbb{P}(1,a_{1},a_{2},a_{3});$
$P\mapsto[g_{0}(P):g_{1}(P):g_{2}(P):g_{3}(P)];$
$\Psi_{a_{4}}:{}$ $Y\dashrightarrow\mathbb{P}(1,a_{1},a_{2},a_{3},a_{4});$
$P\mapsto[g_{0}(P):\dots:g_{4}(P)].$
We claim that they have the following geometric properties.
Claim 7.5.
  1. (1)
    $\Psi_{a_{2}}$ is dominant; $\Psi_{a_{3}}$ is dominant and generically finite of degree $2$;
  2. (2)
    $\Psi_{a_{4}}$ is birational onto its image;
  3. (3)
    let $Y^{\prime}$ be the closure of $\Psi_{a_{4}}(Y)$ in $\mathbb{P}(1,a_{1},a_{2},a_{3},a_{4})$, then $Y^{\prime}$ is defined by a weighted homogeneous polynomial $f$ of degree $d$, where
    $f(x_{0},\dots,x_{4})=x_{4}^{2}+f_{0}(x_{0},x_{1},x_{2},x_{3})$
    in suitable homogeneous coordinates $[x_{0}:\dots:x_{4}]$ of $\mathbb{P}(1,a_{1},a_{2},a_{3},a_{4})$.
Proof.
(1) By Lemma 7.3, $\{g_{0},g_{1},g_{2},g_{3}\}$ is algebraically independent, so $\Psi_{a_{2}}$ and $\Psi_{a_{3}}$ are dominant. In particular, $\Psi_{a_{3}}$ is generically finite by dimension reason. The degree of $\Psi_{a_{3}}$ is the number of points in the fiber over a general point in $\mathbb{P}(1,a_{1},a_{2},a_{3})$. After taking a resolution $\pi:W\to Y$ such that $N_{k}=\operatorname{Mov}|\lfloor{\pi^{*}(a_{k}H)}\rfloor|$ is free for $k=1,2,3$, the degree is just the intersection number $(N_{1}\cdot N_{2}\cdot N_{3})$. So
$\deg(\Psi_{a_{3}})=(N_{1}\cdot N_{2}\cdot N_{3})\leq a_{1}a_{2}a_{3}H^{3}=2.$
On the other hand, it is clear that $\deg(\Psi_{a_{3}})>1$ by the assumption that either $|a_{3}H|$ defines a generically finite map of degree greater than $1$ or $Y$ is not rational.
(2) By assumption, $|a_{4}H|$ defines a birational map. As $g_{4}$ is general, it can separate two points in a general fiber of $\Psi_{a_{3}}$, so $\Psi_{a_{4}}$ is birational onto its image.
(3) Note that $d=2a_{4}$ and $h^{0}(Y,dH)=|S_{2a_{4}}|+|S_{a_{4}}|$. On the other hand,
$S_{2a_{4}}\sqcup(S_{a_{4}}\cdot g_{4})\sqcup\{g_{4}^{2}\}\subseteq H^{0}(Y,dH).$
So $S_{2a_{4}}\sqcup(S_{a_{4}}\cdot g_{4})\sqcup\{g_{4}^{2}\}$ is linearly dependent in $H^{0}(Y,dH)$. In other words, there exists a weighted homogeneous polynomial $f(x_{0},\dots,x_{4})$ of degree $d$ with $\text{wt}(x_{0},\dots,x_{4})=(1,a_{1},a_{2},a_{3},a_{4})$ such that $f(g_{0},\dots,g_{4})=0.$ So $Y^{\prime}$ is contained in $(f=0)\subseteq\mathbb{P}(1,a_{1},a_{2},a_{3},a_{4})$. Note that $Y^{\prime}$ is a hypersurface in $\mathbb{P}(1,a_{1},a_{2},a_{3},a_{4})$ by dimension reason.
We claim that $Y^{\prime}=(f=0)$ and $x_{4}^{2}$ has nonzero coefficient in $f$. Otherwise, either $Y^{\prime}$ is defined by a weighted homogeneous polynomial of degree at most $d-1$, or $x_{4}^{2}$ has zero coefficient in $f$. In either case, $Y^{\prime}$ is defined by a weighted homogeneous polynomial $\tilde{f}$ of the form
$\tilde{f}(x_{0},\dots,x_{4})=x_{4}\tilde{f}_{1}(x_{0},x_{1},x_{2},x_{3})+\tilde{f}_{2}(x_{0},x_{1},x_{2},x_{3}).$
Here $\tilde{f}_{1}\neq 0$ as $\{g_{0},g_{1},g_{2},g_{3}\}$ is algebraically independent. Then $Y^{\prime}$ is birational to $\mathbb{P}(1,a_{1},a_{2},a_{3})$ under the rational projection map
$\mathbb{P}(1,a_{1},a_{2},a_{3},a_{4}){}$ $\dashrightarrow\mathbb{P}(1,a_{1},a_{2},a_{3});$
$[x_{0}:\dots:x_{4}]{}$ $\mapsto[x_{0}:\dots:x_{3}].$
But this contradicts the fact that $\deg(\Psi_{a_{3}})=2$. So $Y^{\prime}=(f=0)$ and $x_{4}^{2}$ has nonzero coefficient in $f$. After a suitable coordinate transformation, we may assume that $f=x_{4}^{2}+f_{0}(x_{0},x_{1},x_{2},x_{3})$. ∎
Now go back to the proof of Theorem 7.1. By Claim 7.5, $f$ is the only algebraic relation on $g_{0},\dots,g_{4}$. Denote $\mathcal{R}$ to be the graded $-\mathbb{C}$-algebra of $R(Y,H)$ generated by $\{g_{0},\dots,g_{4}\}$. Then there is a natural isomorphism between graded $\mathbb{C}$-algebras
$\mathcal{R}\simeq\mathbb{C}[x_{0},\dots,x_{4}]/(x_{4}^{2}+f_{0}(x_{0},x_{1},x_{2},x_{3}))$
by sending $g_{i}\mapsto x_{i}$ ($0\leq i\leq 4$) and the right-hand side is exactly the coordinate ring of $Y^{\prime}$. Write $\mathcal{R}=\bigoplus_{k\geq 0}\mathcal{R}_{k}$, where $\mathcal{R}_{k}$ is the homogeneous part of degree $k$. Then, by [13, 3.4.2],
$\sum_{k\geq 0}(\dim_{\mathbb{C}}\mathcal{R}_{k})q^{k}=\frac{1-q^{d}}{(1-q)(1-q^{a_{1}})(1-q^{a_{2}})(1-q^{a_{3}})(1-q^{a_{4}})}.$
So $\dim_{\mathbb{C}}\mathcal{R}_{k}=h^{0}(X_{d},\mathcal{O}_{X_{d}}(k))=h^{0}(Y,kH)$ for any $k\in\mathbb{Z}_{\geq 0}$, and hence the inclusion $\mathcal{R}\subseteq R(Y,H)$ is an identity. This implies that
$Y\simeq\operatorname{Proj}R(Y,H)=\operatorname{Proj}\mathcal{R}\simeq Y^{\prime}.$
This finishes the proof. ∎
8. Proofs of Theorem 1.1, Theorem 1.2 and Theorem 1.3
8.1. Plurigenera
Proof of Theorems 1.1(1), 1.2(1), and 1.3(1).
These are directly from Propositions 4.7, 5.5, 6.1, and 6.4. ∎
8.2. Weighted embedding
Proof of Theorems 1.1(2) and 1.2(2).
We may replace $W$ by its canonical model $Y$ and apply Theorem 7.1 to $H=K_{Y}$. Then Theorem 1.1(2) is a direct consequence of Propositions 4.7 and 4.8; Theorem 1.2(2) is a direct consequence of Propositions 5.5 and 5.6. ∎
Proof of Theorem 1.3(2).
We may replace $W$ by its canonical model $Y$.
Suppose that the canonical map of $Y$ is generically finite, then we get (2a) by Proposition 6.1.
Suppose that the canonical map of $Y$ is not generically finite. Then we are going to embed $Y$ into $\mathbb{P}(1,1,1,1,2,5)$. Theorem 7.1 is not directly applicable here as we need $2$ defining equations, but the proof is similar to that of Theorem 7.1 with some essential modification.
Since $p_{g}(Y)=4$, we may take $\{g_{0},g_{1},g_{2},g_{3}\}$ as a basis of $H^{0}(Y,K_{Y})$. Take general elements $g_{4}\in H^{0}(Y,2K_{Y})$ and $g_{5}\in H^{0}(Y,5K_{Y})$. We regard $g_{i}$ as elements in $R(Y)=\bigoplus_{k\geq 0}H^{0}(Y,kK_{Y})$.
By Lemma 6.3, the canonical image of $Y$ in $\mathbb{P}^{3}$ is a surface of degree $2$, so $g_{0},g_{1},g_{2},g_{3}$ satisfy a unique quadratic algebraic relation $q(g_{0},g_{1},g_{2},g_{3})=0$. Without loss of generality, we may assume that $q$ is of the form
$q(x_{0},x_{1},x_{2},x_{3})=x_{3}^{2}+q_{0}(x_{0},x_{1},x_{2}).$
In particular, $\{g_{0},g_{1},g_{2}\}$ is algebraically independent and $\mathbb{C}(g_{0},\dots,g_{3})$ is a field extension of $\mathbb{C}(g_{0},g_{1},g_{2})$ of degree $2$. Then, for any $k\geq 0$,
$S_{k}=\{g_{0}^{s_{0}}g_{1}^{s_{1}}g_{2}^{s_{2}}g_{3}^{s_{3}}\mid s_{0},\dots,s_{3}\in\mathbb{Z}_{\geq 0},s_{0}+s_{1}+s_{2}+s_{3}=k\text{ and }s_{3}\leq 1\}$
is linearly independent in $H^{0}(Y,kK_{Y})$. By Proposition 6.4(3),
$H^{0}(Y,2K_{Y})=10=|S_{2}|+1,$
so $S_{2}\sqcup\{g_{4}\}$ forms a basis of $H^{0}(Y,2K_{Y})$. Since $|2K_{Y}|$ defines a generically finite map of degree $2$ by Proposition 6.4(4), $\{g_{0},g_{1},g_{2},g_{4}\}$ is algebraically independent. Then $S_{5}\sqcup(S_{3}\cdot g_{4})\sqcup(S_{1}\cdot g_{4}^{2})$ is linearly independent in $H^{0}(Y,5K_{Y})$. By Proposition 6.4(3),
$H^{0}(Y,5K_{Y})=57=|S_{5}|+|S_{3}|+|S_{1}|+1,$
hence $S_{5}\sqcup(S_{3}\cdot g_{4})\sqcup(S_{1}\cdot g_{4}^{2})\sqcup\{g_{5}\}$ forms a basis of $H^{0}(Y,5K_{Y})$.
Now we can consider $2$ rational maps:
$\Psi_{2}:{}$ $Y\dashrightarrow\mathbb{P}(1,1,1,1,2);$
$P\mapsto[g_{0}(P):\dots:g_{4}(P)];$
$\Psi_{5}:{}$ $Y\dashrightarrow\mathbb{P}(1,1,1,1,2,5);$
$P\mapsto[g_{0}(P):\dots:g_{5}(P)].$
Then $\Psi_{2}$ is generically finite of degree $2$ as $|2K_{Y}|$ defines a generically finite map of degree $2$; moreover, as $|5K_{Y}|$ defines a birational map by Proposition 6.4(5), $g_{5}$ separates fibers of $\Psi_{2}$ and hence $\Psi_{5}$ is birational onto its image.
By Proposition 6.4(3),
$h^{0}(Y,10K_{Y})=342=\sum_{i=0}^{5}|S_{10-2i}|+\sum_{j=0}^{2}|S_{5-2j}|.$
On the other hand,
$\bigsqcup_{i=0}^{5}(S_{10-2i}\cdot g_{4}^{i})\sqcup\bigsqcup_{j=0}^{2}(S_{5-2j}\cdot g_{4}^{j}g_{5})\sqcup\{g_{5}^{2}\}\subseteq H^{0}(Y,10K_{Y}).$
So there is a weighted homogeneous polynomial $f$ of degree $10$ such that $f(g_{0},\dots,g_{5})=0$.
Let $Y^{\prime}$ be the closure of $\Psi_{5}(Y)$ in $\mathbb{P}(1,1,1,1,2,5)$. Then clearly $Y^{\prime}\subseteq(q=f=0)$. We claim that $Y^{\prime}=(q=f=0)$ and $x_{5}^{2}$ has nonzero coefficient in $f$. Otherwise, $Y^{\prime}$ is contained in $(q=\tilde{f}=0)$, where $\tilde{f}$ of the form
$\tilde{f}(x_{0},\dots,x_{5})=x_{5}\tilde{f}_{1}(x_{0},\dots,x_{4})+\tilde{f}_{2}(x_{0},\dots,x_{4}).$
Here $\tilde{f}$ is not divisible by $q$, which implies that $\tilde{f}_{1}\neq 0$. But then $Y^{\prime}$ is birational to $\Psi_{2}(Y)$ under the rational projection map
$\mathbb{P}(1,1,1,1,2,5)\dashrightarrow\mathbb{P}(1,1,1,1,2),$
which contradicts the fact that $\Psi_{2}$ is not birational. So $Y^{\prime}=(q=f=0)$ and $x_{5}^{2}$ has nonzero coefficient in $f$. After a suitable coordinate transformation, we may assume that $f=x_{5}^{2}+f_{0}(x_{0},\dots,x_{4})$.
Now $q$ and $f$ are the only algebraic relations on $g_{0},\dots,g_{5}$. Denote $\mathcal{R}$ to be the graded $-\mathbb{C}$-algebra of $R(Y)$ generated by $\{g_{0},\dots,g_{5}\}$. Then there is a natural isomorphism between graded $\mathbb{C}$-algebras
$\mathcal{R}\simeq\mathbb{C}[x_{0},\dots,x_{5}]/(q,f)$
by sending $g_{i}\mapsto x_{i}$ ($0\leq i\leq 5$) and the right-hand side is exactly the coordinate ring of $Y^{\prime}$. Write $\mathcal{R}=\bigoplus_{k\geq 0}\mathcal{R}_{k}$, where $\mathcal{R}_{k}$ is the homogeneous part of degree $k$. Then, by [13, 3.4.2],
$\sum_{k\geq 0}(\dim_{\mathbb{C}}\mathcal{R}_{k})q^{k}=\frac{(1-q^{2})(1-q^{10})}{(1-q)^{4}(1-q^{2})(1-q^{5})}=\frac{1-q^{10}}{(1-q)^{4}(1-q^{5})}.$
So $\dim_{\mathbb{C}}\mathcal{R}_{k}=h^{0}(Y,kK_{Y})$ for any $k\in\mathbb{Z}_{\geq 0}$ by Proposition 6.4(3), and hence the inclusion $\mathcal{R}\subseteq R(Y)$ is an identity. This implies that
$Y\simeq\operatorname{Proj}R(Y)=\operatorname{Proj}\mathcal{R}\simeq Y^{\prime}.$
This finishes the proof of (2b).
Finally, for a $3$-fold $Y$ defined by $(q=f=0)$ in $\mathbb{P}(1,1,1,1,2,5)$, we can view it as a specialization of $Y_{t}$ ($t\in\mathbb{C}^{*}$) which is defined by
$(q+tx_{4}=f+tg=0),$
where $g$ is a general weighted polynomial in $x_{0},\dots,x_{3},x_{5}$ of degree $10$. In particular, $(g=0)\subseteq\mathbb{P}(1,1,1,1,5)$ is smooth. Then $Y_{t}$ turns out to be a $3$-fold defined by
$f(x_{0},x_{1},x_{2},x_{3},-\frac{1}{t}q,x_{5})+tg(x_{0},x_{1},x_{2},x_{3},x_{5})=0$
in $\mathbb{P}(1,1,1,1,5)$ with coordinates $[x_{0}:x_{1}:x_{2}:x_{3}:x_{5}]$, which is actually smooth for general $t\neq 0$. ∎
8.3. Simply connectedness
Proof of Theorems 1.1(3), 1.2(3), and 1.3(3).
Let $W$ be as in the assumption. By (2) of each theorem, the canonical model $Y$ of $W$ is a specialization of general hypersurfaces in a certain weighted projective space. Namely, there is a flat morphism $\mathcal{Y}\to C$ to a smooth curve such that the central fiber $\mathcal{Y}_{o}$ at $o\in C$ is isomorphic to $Y$ and a general fiber $\mathcal{Y}_{t}$ for $t\neq o$ is a general weighted hypersurface. By [23, Lemma 2.8.1], after shrinking $C$, we may assume that $\mathcal{Y}\setminus\mathcal{Y}_{o}\to C\setminus\{o\}$ is a topological fiber bundle. Note that $\mathcal{Y}$ is normal by [27, Theorem 23.9]. Hence by [23, Theorem 2.12], there is a surjective map $\pi_{1}(\mathcal{Y}_{t})\to\pi_{1}(\mathcal{Y}_{o})$ for general $t\in C$. Such $\mathcal{Y}_{t}$ is quasismooth by [18] and hence is simply connected by [13, Theorem 3.2.4]. Hence $Y$ is also simply connected. Since fundamental group is a birational invariant among projective klt varieties by [32, Theorem 1.1], we conclude the simply connectedness of $W$. ∎
8.4. Irreducibility and dimensions of moduli spaces
Proof of Theorem 1.1(3).
Write $\mathbb{P}=\mathbb{P}(1,1,2,3,8)$. By Theorem 1.1(2), it is clear that $\mathcal{M}_{\frac{1}{3},2}$ is irreducible. It is unirational as it is dominated by an open subvariety of a rational variety (parametrized by coefficients). Furthermore,
$\dim\mathcal{M}_{\frac{1}{3},2}=h^{0}(\mathbb{P},\mathcal{O}_{\mathbb{P}}(16))-1-\dim\operatorname{Aut}\mathbb{P}.$
By an easy computation, $h^{0}(\mathbb{P},\mathcal{O}_{\mathbb{P}}(16))=246$. By [11, §4] and [12, Proposition 4.3], we have
$\dim\operatorname{Aut}(\mathbb{P})$ $=2h^{0}(\mathcal{O}_{\mathbb{P}}(1))+h^{0}(\mathcal{O}_{\mathbb{P}}(2))+h^{0}(\mathcal{O}_{\mathbb{P}}(3))+h^{0}(\mathcal{O}_{\mathbb{P}}(8))-1$
$=56.$
Therefore, $\dim\mathcal{M}_{\frac{1}{3},2}=189$. ∎
Proof of Theorem 1.2(3).
Write $\mathbb{P}=\mathbb{P}(1,1,1,2,6)$. By Theorem 1.2(2), it is clear that $\mathcal{M}_{1,3}$ is irreducible. It is unirational as it is dominated by an open subvariety of a rational variety (parametrized by coefficients). Furthermore,
$\dim\mathcal{M}_{1,3}=h^{0}(\mathbb{P},\mathcal{O}_{\mathbb{P}}(12))-1-\dim\operatorname{Aut}\mathbb{P}.$
By an easy computation, $h^{0}(\mathbb{P},\mathcal{O}_{\mathbb{P}}(12))=303$. By [11, §4] and [12, Proposition 4.3], we have
$\dim\operatorname{Aut}(\mathbb{P})$ $=3h^{0}(\mathcal{O}_{\mathbb{P}}(1))+h^{0}(\mathcal{O}_{\mathbb{P}}(2))+h^{0}(\mathcal{O}_{\mathbb{P}}(6))-1=66.$
Therefore, $\dim\mathcal{M}_{1,3}=236$. ∎
Proof of Theorem 1.3(4).
Write $\mathbb{P}=\mathbb{P}(1,1,1,1,5)$. By Theorem 1.3(2), it is clear that $\mathcal{M}_{2,4}$ is irreducible. It is unirational as it is dominated by an open subvariety of a rational variety (parametrized by coefficients). Furthermore,
$\dim\mathcal{M}_{2,4}=h^{0}(\mathbb{P},\mathcal{O}_{\mathbb{P}}(10))-1-\dim\operatorname{Aut}\mathbb{P}.$
By an easy computation, $h^{0}(\mathbb{P},\mathcal{O}_{\mathbb{P}}(10))=343$. By [11, §4] and [12, Proposition 4.3], we have
$\dim\operatorname{Aut}(\mathbb{P})$ $=4h^{0}(\mathcal{O}_{\mathbb{P}}(1))+h^{0}(\mathcal{O}_{\mathbb{P}}(5))-1=72.$
Therefore, $\dim\mathcal{M}_{2,4}=270$. ∎
9. Lower bounds for canonical volumes of $3$-folds of general type with $5\leq p_{g}\leq 10$
Throughout this section, we always assume that $X$ is a minimal $3$-fold of general type with $p_{g}(X)\geq 5$.
Setting 9.1.
Let $\pi:W\to X$ be as in §2.3 with respect to $|K_{X}|$ and keep the same notation. We have the following commutative diagram:
${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{\gamma}$$\scriptstyle{\psi}$${\Sigma^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{|K_{X}|}}$${\Sigma}$
We may write
$K_{W}{}$ $=\pi^{*}(K_{X})+E_{\pi},$
$\pi^{*}(K_{X}){}$ $=M+Z_{\pi},$
where $|M|=\operatorname{Mov}|\lfloor{\pi^{*}(K_{X})}\rfloor|$, $E_{\pi},Z_{\pi}$ are effective $\mathbb{Q}$-divisor.
Let $S$ be a general irreducible element of $|M|$. Then $S$ is a smooth projective surface of general type. Denote by $\sigma_{S}:S\to S_{0}$ the contraction to its minimal model.
After taking some further blow-up, we may further assume that the following hold:
  1. (1)
    $\operatorname{Mov}|K_{S}|$ is base point free;
  2. (2)
    $\operatorname{Supp}(Z_{\pi}+E_{\pi})$ is of simple normal crossing;
  3. (3)
    $\pi$ factorizes through the modification $\mu:X^{\prime}\to X$ in Setting 3.2.
We first consider the case when $|K_{X}|$ is not composed with a pencil of $(1,2)$-surfaces.
Proposition 9.2.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)\geq 5$. Suppose that $|K_{X}|$ is not composed with a pencil of $(1,2)$-surfaces. Then
$K_{X}^{3}\geq\frac{1}{4}\lceil{\frac{8}{3}(2p_{g}(X)-5)}\rceil.$
Proof.
Keep the notation in Setting 9.1.
If $\dim\Sigma=3$, then by [22, Theorem 2.4], we have $K_{X}^{3}\geq 2p_{g}(X)-6.$
If $\dim\Sigma=1$, since $|K_{X}|$ is not composed with a pencil of $(1,2)$-surfaces, then by [4, Theorem 4.4, Theorem 4.5], we have $K_{X}^{3}>2p_{g}(X)-6.$
If $\dim\Sigma=2$ and a general fiber $C$ of $\psi$ is of genus at least $3$, then by [4, Theorem 4.1], we have $K_{X}^{3}\geq 2p_{g}(X)-4$.
From now on, suppose that $\dim\Sigma=2$ and a general fiber $C$ of $\psi$ is of genus $2$. Since $\dim\Sigma=2$, we may write $S|_{S}\equiv dC$, where $d=\deg(s)\deg(\Sigma)\geq p_{g}(X)-2$. By the adjunction formula,
$K_{S}$ $=(K_{X^{\prime}}+S)|_{S}=(\pi^{*}(K_{X})+S+E_{\pi})|_{S}$
$=2S|_{S}+(E_{\pi}+Z_{\pi})|_{S}\equiv 2dC+(E_{\pi}+Z_{\pi})|_{S}.$
It is clear that $p_{g}(S)\geq p_{g}(X)-1\geq 3$. By [4, Proposition 2.9], $K_{S_{0}}^{2}\geq\frac{8}{3}(2d-1)$. Since $K_{S_{0}}^{2}$ is an integer, we deduce that
$K_{S_{0}}^{2}\geq\lceil{\frac{8}{3}(2p_{g}(X)-5)}\rceil.$
By [4, Corollary 2.3] (take $D=K_{X}$ and $\lambda=1$), $\pi^{*}(K_{X})|_{S}-\frac{1}{2}\sigma_{S}^{*}(K_{S_{0}})$ is $\mathbb{Q}$-effective. Hence
$K_{X}^{3}\geq(\pi^{*}(K_{X})|_{S})^{2}\geq\frac{1}{4}K_{S_{0}}^{2}\geq\frac{1}{4}\lceil{\frac{8}{3}(2p_{g}(X)-5)}\rceil.$
The proof is completed. ∎
From now on, we consider the case that $|K_{X}|$ is composed with a pencil of $(1,2)$-surfaces.
Proposition 9.3.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)\geq 5$. Suppose that $|K_{X}|$ is composed with a pencil which is not rational. Then $K_{X}^{3}\geq p_{g}(X)$.
Proof.
Keep the notation in Setting 9.1. Then we may write
$\pi^{*}(K_{X})\equiv dS+Z_{\pi},$
where $d\geq p_{g}(X)$ as the pencil is not rational. Then by [3, Lemma 2.1(i)],
$(\pi^{*}(K_{X})|_{S})^{2}=\mathrm{Vol}(S)\geq 1.$
Then
$K_{X}^{3}\geq d(\pi^{*}(K_{X})|_{S})^{2}\geq p_{g}(X).$
The proof is completed. ∎
The following is the key proposition of this section, which deals with the most difficult case.
Proposition 9.4.
Let $X$ be a minimal $3$-fold of general type with $p_{g}(X)\geq 5$. Suppose that $|K_{X}|$ is composed with a rational pencil of $(1,2)$-surfaces. Then
$K_{X}^{3}\geq\min\left\{\frac{n_{0}}{m_{0}}\left(\frac{p_{g}(X)-1}{p_{g}(X)}+\frac{m_{0}-2}{2}\right)+\frac{(p_{g}(X)-1)^{2}}{p_{g}(X)},\right.$
$\left.\max\left\{p_{g}(X)-1,\frac{4}{3}p_{g}(X)-\frac{10}{3}\right\}\right\}$
where $m_{0}=\lfloor{\frac{7p_{g}(X)-1}{10}}\rfloor$, $n_{0}=\lfloor{\frac{7p_{g}(X)-21}{10}}\rfloor$.
Proof.
Keep the notation in Setting 9.1. We may write
$\pi^{*}(K_{X})=(p_{g}-1)S+Z_{\pi}.$
Here we simply write $p_{g}=p_{g}(X)$.
Take integers $m\geq 2$ and $n\geq 0$ satisfying
$\left(m-1-\frac{n+1}{p_{g}-1}\right)\cdot\frac{p_{g}-1}{p_{g}}>m-2+\frac{3}{10},$ (9.1)
or equivalently,
$m+n<\frac{7p_{g}}{10}.$
Note that
$(m-1)\pi^{*}(K_{X})-(n+1)S-\frac{n+1}{p_{g}-1}Z_{\pi}\equiv\left(m-1-\frac{n+1}{p_{g}-1}\right)\pi^{*}(K_{X})$
is nef and big. By the Kawamata–Viehweg vanishing theorem,
$H^{1}(W,K_{W}+\lceil{(m-1)\pi^{*}(K_{X})-(n+1)S-\frac{n+1}{p_{g}-1}Z_{\pi}}\rceil)=0,$
which implies that the natural restriction map
$H^{0}(W,K_{W}+\lceil{(m-1)\pi^{*}(K_{X})-nS-\frac{n+1}{p_{g}-1}Z_{\pi}}\rceil)$
[…]
is surjective. Here in the last term, the restriction on $S$ commutes with the roundup as $S$ is general, and we use the fact that $S|_{S}\sim 0$. Therefore,
$|mK_{W}-nS||_{S}\succcurlyeq|K_{S}+\lceil{L_{m,n}}\rceil+(m-2)\sigma_{S}^{*}(K_{S_{0}})|,$ (9.2)
where
$L_{m,n}=(m-1)\pi^{*}(K_{X})|_{S}-\frac{n+1}{p_{g}-1}Z_{\pi}|_{S}-(m-2)\sigma_{S}^{*}(K_{S_{0}}).$
Recall that by our construction, $\pi$ factorizes through $\mu$ in Setting 3.2. Hence by Lemma 3.3,
$L_{m,n}\geq{}$ $\left(m-1-\frac{n+1}{p_{g}-1}\right)\pi^{*}(K_{X})|_{S}-(m-2)\sigma_{S}^{*}(K_{S_{0}})$
[…]
In the last step, the coefficient of $\sigma_{S}^{*}(K_{S_{0}})$ is strictly larger than $\frac{3}{10}$ by (9.1). Recall that $S$ is a $(1,2)$-surface by assumption, so by [2, Theorem 3.2], $h^{0}(S,K_{S}+\lceil{L_{m,n}}\rceil)\geq 3$.
Denote $|H_{m,n}|=\operatorname{Mov}|K_{S}+\lceil{L_{m,n}}\rceil|$, then by comparing the movable parts of both sides of (9.2), we have
$M_{m,-n}|_{S}\geq H_{m,n}+(m-2)C,$ (9.3)
where $|M_{m,-n}|=\operatorname{Mov}|\lfloor{\pi^{*}(mK_{X})-nS}\rfloor|$ and $C\in\operatorname{Mov}|K_{S}|=\operatorname{Mov}|\sigma_{S}^{*}(K_{S_{0}})|$ is a general element. Here we recall that
$\operatorname{Mov}|\lfloor{\pi^{*}(mK_{X})-nS}\rfloor|=\operatorname{Mov}|mK_{W}-nS|$
by the isomorphism $H^{0}(W,mK_{W})\simeq H^{0}(W,\lfloor{\pi^{*}(mK_{X})}\rfloor)$.
Recall that by [7, §3.7],
$(\pi^{*}(K_{X})\cdot C)=\xi\geq 1.$ (9.4)
If $|H_{m,n}|$ is composed with a pencil, then it is composed with the same pencil as $|K_{S}|$ and hence $H_{m,n}\geq 2C$ as $h^{0}(S,H_{m,n})\geq 3$. Then (9.3) implies that $\pi^{*}(mK_{X})|_{S}\geq M_{m,-n}|_{S}\geq mC$, so $\pi^{*}(K_{X})|_{S}-C$ is $\mathbb{Q}$-effective. In this case, by (9.4),
$K_{X}^{3}\geq(p_{g}-1)(\pi^{*}(K_{X})|_{S})^{2}\geq(p_{g}-1)(\pi^{*}(K_{X})\cdot C)\geq p_{g}-1.$
Also we have $(\pi^{*}(K_{X})|_{S})^{2}\geq(\pi^{*}(K_{X})|_{S}\cdot C)\geq 1$. By $K_{S}\geq\pi^{*}(K_{X})|_{S}$ and the Zariski decomposition, we have $\sigma_{S}^{*}(K_{S_{0}})\geq\pi^{*}(K_{X})|_{S}$. By assumption, $K_{S_{0}}^{2}=1=(\pi^{*}(K_{X})|_{S})^{2}$. So by Lemma 2.2, $\sigma_{S}^{*}(K_{S_{0}})=\pi^{*}(K_{X})|_{S}$. Then by [4, Lemma 3.2] and [4, Proof of Corollary 3.3(3)], after replacing $X$ by another minimal model, we may assume that $\text{\rm Mov}|K_{X}|$ is base point free. Then by [4, Theorem 4.7], $K_{X}^{3}\geq\frac{4}{3}p_{g}-\frac{10}{3}$.
So in the following, we suppose that $|H_{m,n}|$ is not composed with a pencil for any $m,n$ satisfying (9.1). In this case, $(H_{m,n}\cdot C)\geq 2$. On the other hand, by Lemma 3.3, $\pi^{*}(K_{X})|_{S}-\frac{p_{g}-1}{p_{g}}C$ is $\mathbb{Q}$-effective. Hence
$(\pi^{*}(K_{X})|_{S}\cdot H_{m,n})\geq\frac{p_{g}-1}{p_{g}}(C\cdot H_{m,n})\geq\frac{2p_{g}-2}{p_{g}}.$ (9.5)
Then we consider special values of $(m,n)$. Take $m_{0}=\lfloor{\frac{7p_{g}(X)-1}{10}}\rfloor$ and $n_{0}=\lfloor{\frac{7p_{g}(X)-21}{10}}\rfloor$.
Note that $(m,n)=(m_{0},0)$ satisfies (9.1). Then by (9.3),
$\pi^{*}(m_{0}K_{X})|_{S}\geq M_{m_{0},0}|_{S}\geq H_{m_{0},0}+(m_{0}-2)C.$
By (9.4) and (9.5),
$m_{0}(\pi^{*}(K_{X})|_{S})^{2}\geq{}$ $(\pi^{*}(K_{X})|_{S}\cdot H_{m_{0},0})+(m_{0}-2)(\pi^{*}(K_{X})|_{S}\cdot C)$
[…]
On the other hand, $(m,n)=(2,n_{0})$ satisfies (9.1). So by (9.3), $M_{2,-n_{0}}|_{S}\geq H_{2,n_{0}}$. By definition, $\pi^{*}(2K_{X})\geq n_{0}S+M_{2,-n_{0}}$. So
$2K_{X}^{3}\geq{}$ $(\pi^{*}(K_{X})^{2}\cdot(n_{0}S+M_{2,-n_{0}}))$
[…]
Here for the last step we use (9.5) and (9.6). The proof is completed. ∎
By Propositions 9.2, 9.3, and 9.4, we have the following theorem.
Theorem 9.5.
Let $X$ be a minimal $3$-fold of general type with $5\leq p_{g}(X)\leq 10$. Then the lower bound of $K_{X}^{3}$ is given by the following table:
$p_{g}(X)=$ $5$ $6$ $7$ $8$ $9$ $10$
$K_{X}^{3}\geq$ $7/2$ $19/4$ $6$ $22/3$ $26/3$ $10$
Remark 9.6.
  1. (1)
    When $p_{g}(X)=5$, the lower bound in Theorem 9.5 is sharp by an example [9, Table 10, No. 2] and this example justifies the optimality of Proposition 9.2.
  2. (2)
    More surprisingly, if we only look at the first term in the lower bound given by Proposition 9.4, then it is $\frac{109}{30},\frac{85}{14},\frac{151}{20},\frac{301}{30}$ respectively when $p_{g}(X)=5,7,8,10$. Such values are attained by examples [9, Table 10, No. 3, No. 8, No. 9, No. 10].
  3. (3)
    If we apply methods in [17], then we can improve the lower bounds of Theorem 9.5 when $p_{g}(X)=6,8,9$, but we will not write down the details here.
Acknowledgments
This work was supported by National Key Research and Development Program of China #2023YFA1010600, #2020YFA0713200, and NSFC for Innovative Research Groups #12121001. Chen was supported by NSFC grant #12071078. Hu was supported by NSFC grant #12201397. Chen and Jiang are members of the Key Laboratory of Mathematics for Nonlinear Sciences, Fudan University.
We are grateful to Jungkai A. Chen, Stephen Coughlan, Roberto Pignatelli, Mingchen Xia and Tong Zhang for discussions.